### v1

# Identifying treatment response subgroups in observational time-to-event data

Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for subgroup analysis primarily focus on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised. Furthermore, the patient cohort of an RCT is often constrained by cost, and is not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice. Therefore, when applied to observational studies, such approaches suffer from significant statistical biases because of the non-randomisation of treatment. Our work introduces a novel, outcome-guided method for identifying treatment response subgroups in observational studies. Our approach assigns each patient to a subgroup associated with two time-to-event distributions: one under treatment and one under control regime. It hence positions itself in between individualised and average treatment effect estimation. The assumptions of our model result in a simple correction of the statistical bias from treatment non-randomisation through inverse propensity weighting. In experiments, our approach significantly outperforms the current state-of-the-art method for outcome-guided subgroup analysis in both randomised and observational treatment regimes.
Data and Code Availability
Experiments are performed on a synthetic dataset and a real-world dataset from the Surveillance, Epidemiology, and End Results program [^1] Available at https://seer.cancer.gov/. For reproducibility, implementation of the method, synthetic dataset, and analyses are available on Github  [^2] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
1 Introduction
Understanding heterogeneous therapeutic responses between patient subgroups is the core of treatment guidelines and the development of new drugs. Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed [11]. For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes [18], shaping subsequent guidelines. Figure 1 illustrates this idea: two groups may present opposite treatment responses, and would benefit from different recommendations. Our work aims to uncover such subgroups.
Randomised controlled trials (RCTs) remain the gold standard for identifying subgroups of treatment effects. In RCTs, patients are randomly assigned to control or treated groups, allowing researchers to assess the impact of an intervention or treatment. However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies [16].
Our work introduces a novel approach that diverges from traditional RCT-based methodologies [34, 36] by leveraging routinely collected observational data to identify patient subgroups. Observational studies encompass larger and more diverse cohorts reflective of real-world practices, offering the potential to uncover subgroups of treatment responses, that could be missed in RCTs. Prior works [5, 9, 29] have leveraged observational data while addressing biases from non-random treatment assignments [4, 15, 14]. However, this body of literature primarily focuses on estimating (i) averaged treatment effects at the population level or (ii) individualised treatment effects, thereby overlooking the identification of treatment effect subgroups.
Addressing this critical gap in the literature, our work uncovers patient subgroups with distinct treatment responses using observational data. Unlike previous methods relying on RCTs, we introduce a mixture of neural networks as an extension to traditional outcome-guided models to uncover subgroups delineated by non-linear, higher-dimensional combinations of available covariates, without requiring parametrisation of the time-to-event distributions or treatment responses.
Figure 1: Subgroup treatment effect discovery in time-to-event observational data. Our method identifies subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials. Our method simultaneously models the treatment effect and identifies subgroup while addressing censoring and treatment non-randomisation.
Contributions.
Our work extends Neural Survival Clustering [19] to estimate treatment effects, introducing the novel Causal Neural Survival Clustering (CNSC). This methodology is the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects.
Section 2 formalises the problem of treatment subgroup discovery and non-randomisation correction. Then, we introduce the proposed monotonic neural network implementation in Section 2. Finally, we evaluate the proposed methodology on two datasets in Section 3 and compare it with related approaches described in Section 4.
2 Method
In this section, we describe our method, referring to Appendix A for technical proofs.
2.1 Problem setup
Latent survival subgroup analysis.
Our goal is to uncover subgroups with similar treatment responses, guided by the observed times of occurrence of an outcome of interest. Patients are assigned to subgroups based on their covariates at treatment time. Each subgroup is characterised by two distributions, known as survival functions: one under treatment and one under control regimes. These distributions characterise the probability of observing the event of interest after a given time under a given treatment.
Consider the random variables associated with observed covariates $X$, an indicator identifying whether the event of interest, in our analysis death, was observed $D$, and the observed time of event $T$. Formally, we define the latter random variables as $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is random variable of the (right)-censoring time, i.e. the time at which the patient left the study before experiencing the event of interest, and $T^{\prime}$ is the partially observed random variable associated with the time of the event of interest if there was no censoring. When a patient is censored, $T=C$ and $D=0$, otherwise the event is observed and $T=T^{\prime}$ and $D=1$. To estimate the associated survival likelihood, we further assume, as commonly done in survival analysis, that the censoring time $C$ is not informative for the time of the event of interest $T^{\prime}$.
Assumption 1 (Non-informative censoring).
The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$ given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
Central to our problem is the additional variable $Z$ associated with the latent, unobserved subgroup membership. Following Jeanselme et al. [19], we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^3] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3.. We assume that $Z$ depends on $X$ and influences $T^{\prime}$. As we are interested in recovering the survival functions associated with the latent subgroups, we ignore the potential dependence of $T^{\prime}$ on $X$, considering the individual survival distribution as a mixture of the different subgroups.
Assumption 2 (Mixture modelling).
The event time $T^{\prime}$ is completely determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
While this assumption may hurt individual performance as individual covariates do not directly inform individual survival but group membership only, it improves the interpretability of the model by separating subgroup membership from survival profiles. In conclusion, we aim to recover $Z$ and the survival distributions associated with each subgroup from the observed $X$, $T$ and $D$.
Given the previously described dependencies, the expected survival $S$ for a given patient with covariates $x$ is the survival when marginalising over the different subgroups:
$S(t\mid X=x):=\mathbb{P}(T^{\prime}\geq t\mid X=x)=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid X=x)\mathbb{P}(T^{\prime}\geq t\mid Z=k)$ (1)
Note that the last term is linked to the quantity commonly used in the survival literature: $\Lambda_{k}(t):=-\log\mathbb{P}(T^{\prime}\geq t\mid Z=k)$, the group-specific cumulative hazard, with derivative with respect to time $\lambda_{k}(t)$, the group-specific instantaneous hazard, corresponding to the increase in the probability of observing the event of interest given survival until that time $t$. Estimating the two probability distributions $\mathbb{P}(Z=k\mid X=x)$ and $\mathbb{P}(T^{\prime}\geq t\mid Z=k)$ is our core interest.
Likelihood-based model fitting.
One can maximise the log-likelihood associated with the observed realisation $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i$ (we omit $i$ where redundant) to estimate the survival function (Eq. (1)). Under Assumption 1, Equation (2) describes this factual [^4] Referring to the likelihood of the observed (rather than counterfactual) realisations. log-likelihood.
$l_{F}(\theta)$ $=\sum_{i,d_{i}=1}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}\log S_{\theta}(t_{i}\mid x_{i})$ (2)
where $\theta$ is the set of parameters characterising the estimated survival function $S_{\theta}$. Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see [19] for derivation). Each censored patient ($d_{i}=0$) contributes the probability of not observing an event before $t_{i}$, i.e., $S_{\theta}(t_{i}\mid x_{i})$, under Assumption 1 of independence between the censoring time and the event of interest conditional on the covariates.
Latent treatment effect subgroups.
Consider now the binary treatment variable $A$. Patients either receive the treatment ($A=1$) or they do not ($A=0$). Therefore, following the potential outcomes formulation, we consider the time of events under treatment $T^{\prime}_{1}$ and under the control regime $T^{\prime}_{0}$. The central challenge is that one can not observe both $T^{\prime}_{0}$ and $T^{\prime}_{1}$, but only $T^{\prime}$, the random variable associated with the event time under the observed treatment regime in the absence of censoring, which is equal to $A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under Assumption 4 described below. Figure 2 describes the dependencies between the previously described random variables.
A critical assumption of our proposed setting is that the subgroups we aim to identify do not influence treatment if we know a patient’s covariates, formalised as follows:
Assumption 3 (Unknown latent groups).
The treatment assignment $A$ is independent of the subgroup membership $Z$ given the covariates $X$. Formally, $A\perp\!\!\!\perp Z\mid X$.
Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed. $Z$ stands for the latent subgroup structure explaining the event time under no treatment $T^{\prime}_{0}$ and under treatment $T^{\prime}_{1}$. $A$ is the assigned treatment. $C$ is the censoring time.
This assumption is realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) subgroups are unknown and would consequently not inform treatment. If subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
Further, the following three assumptions are necessary to estimate treatment effects and are common in the causal literature.
Assumption 4 (Consistency).
A patient’s observed event time is the potential event time associated with the observed treatment. Formally, this means $T^{\prime}:=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the treatment $A$.
Assumption 5 (Ignorability).
The potential event times are independent of the treatment given the observed covariates, i.e. $A\perp\!\!\!\perp(T^{\prime}_{0},T^{\prime}_{1})\mid X$. Equivalently, no unobserved confounders impact both treatment and event time.
Assumption 6 (Overlap / Positivity).
Each patient has a non-zero probability of receiving the treatment, i.e. $\mathbb{P}(A\mid X)\in(0,1)$ where $(0,1)$ is the open interval, resulting in a non-deterministic treatment assignment.
Under the Assumptions 4, 5, and 6, existing works often focus on estimating the Individualised Treatment Effect (ITE, here denoted as $\tau$), defined as the difference between survival under the two potential treatment regimes given the covariates $x$ (see App. A.1 for derivation):
$\tau(t,x)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (3)
Estimating this quantity requires accurately modelling the survival distributions under the two treatment regimes for all $x$ from observed data. This estimation would be straightforward if one could access the observed survival times for all patients under both treatment regimes. In this case, one would estimate the survival using the event times under $A=1$ and under $A=0$. The key challenge is that the counterfactual survival outcome is not observed: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
In RCTs, treatment $A$ and covariates $X$ are by design independent. Hence, the subset of the patients receiving either treatment regime is representative of the overall population. Relying on this observed subset to estimate the survival distribution under each regime is a valid estimate for the population. In other words, maximising the factual likelihood to estimate survival under both treatment regimes results in a valid estimate. However, covariates and treatment are not independent in observational studies in which treatment recommendations depend upon the observed covariates. Formally, $\mathbb{P}(A\mid X)\neq\mathbb{P}(A)$ in general. For example, clinicians might recommend more aggressive treatment to patients in more severe conditions. This absence of randomisation results in a covariate shift between the treated and non-treated populations [9] as their covariate distributions differ [6]. In this setting, estimating survival by maximising the factual likelihood is no-longer enough to estimate the counterfactual survival distribution.
Under Assumption 6, approaches such as re-weighting [48], or penalisation on the dissimilarity of learnt representations [22, 47], or a combination of these approaches [9, 13, 45] are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest. Specifically, Shalit et al. [45] demonstrate that the negative log likelihood is upper-bounded by observable quantities. Here we extend their notations to the survival setting as
$-l(\theta)\leq-l_{F}^{*}(\theta)+\gamma\cdot\text{IPM}(q_{\Phi}^{A=0},q_{\Phi}^{A=1}),$ (4)
where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in Eq.(2)) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
Shalit et al. propose to use this bound to train a neural network to estimate treatment effects. In this context, $\Phi$ is an inner representation of the network and the IPM regularisation renders this embedding similar between the treated and untreated populations. This embedding is then used to estimate the survival function under both treatment regimes. The difference between the estimated survival functions is then an accurate estimate of the treatment effect as the IPM penalisation corrects for the shift resulting from treatment non-randomisation.
Note that, under the dependencies we assume, only the subgroup assignment ($\mathbb{P}(Z\mid X)$) depends on the covariates $X$ and could, consequently, play the role of the transformation $\Phi$. However, treatment is assumed independent of the group membership under Assumption 3, meaning that treatment rate does not differ between subgroups. This assumption results in the regularisation term being null, i.e. $\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$ (see App. A.2 for derivation).
As a consequence, our work focuses on the first term of the upper bound in Eq. (4) by re-weighting the factual likelihood with the patient’s propensity score [13, 45]. Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting [3] scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
$w_{i}^{-1}=\begin{cases}0.05&\text{ if }\hat{p}_{A}(x_{i})<0.05\\ 0.95&\text{ if }\hat{p}_{A}(x_{i})>0.95\\ a_{i}\cdot\hat{p}_{A}(x_{i})+(1-a_{i})\cdot(1-\hat{p}_{A}(x_{i}))&\text{ otherwise},\end{cases}$ (5)
where $i$ is the patient index, and $a_{i}$ is the realisation of $A$ for patient $i$. Using the factual likelihood from Eq. (2) and the weights $w_{i}$ defined in Eq. (5), we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
$\sum_{i,d_{i}=1}w_{i}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i},a_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}w_{i}\log S_{\theta}(t_{i}\mid x_{i},a_{i})$ (6)
Figure 3: Causal Neural Survival Clustering architecture. Latent parameter $l_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the cumulative hazard $\Lambda_{k}$ under both treatment regimes. $G$ assigns the probability to belong to each subgroup given the patient’s covariate(s) $x$. To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood. In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
2.2 Estimating the quantities of interest with neural networks
The previous section discussed the quantities one must estimate —here parameterised by neural networks— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the cumulative incidence $\Lambda_{k}$ characterising survival under the two treatment regimes, and the weights $w$. Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
Subgroup assignment.
Similar to [19], a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
$G(x):=[\mathbb{P}(Z=k\mid x)]_{k=1}^{K}$
Survival distributions.
Each subgroup $k$ is represented by a vector $l_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the cumulative hazard functions. The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^5] To ensure this constraint, we apply a square function on all weights as proposed in [19, 20]., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$. We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks [46]:
$\Lambda_{k}(t):=t\cdot M(l_{k},t)$
By modelling these two cumulative intensity functions, we can estimate the Group Average Treatment Effect (GATE), $\hat{\tau}_{k}(t)$, in subgroup $k$ as
$\hat{\tau}_{k}(t):=\mathbb{E}[\tau(t,X)\mid Z=k)]=\exp(-\Lambda_{k}(t\mid A=1))-\exp(-\Lambda_{k}(t\mid A=0))$
Inverse propensity weighting.
Under treatment randomisation, as in RCTs, the previous components would accurately model the observed survival outcomes and identify subgroups of treatment effects by maximising the factual likelihood. As previously discussed, to account for the treatment non-randomisation in observational studies, we weigh the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
$W(x):=\mathbb{P}(A=1\mid x)$
2.3 Training procedure
First, the network $W$ is trained to predict the binary treatment assignment by minimising the cross-entropy of receiving treatment. Then, training all other components relies on minimising the weighted factual log-likelihood introduced in Eq. (6). The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative used in the log likelihood [19, 20, 40].
3 Experimental analysis
As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known. Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer. Our code for the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^6] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
3.1 Data generation
We generate a population of $30,000$ divided into $K=3$ subgroups. We draw 10 covariates from normal distributions and survival distributions for each treatment regime following Gompertz distributions [37] parameterised by group membership and individual covariates. This choice reflects a setting in which time-to-event distributions under treatment or control regimes are different functions of the covariates. This simulation is more likely to capture the complexity of real-world responses, contrary to traditional evaluations of subgroup analysis, which often assume a linear treatment response, i.e., a simple shift between treated and untreated distributions. Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as observed in observational studies (Observational). Finally, non-informative censoring times are drawn also following a Gompertz distribution. Details of the generative process of our synthetic dataset are deferred to App. B.1.
3.2 Empirical settings
Benchmarks.
We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE [34] [^7] Implemented in the Auton-Survival library [35]) which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach is that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies subgroups of treatment while considering survival, without constraining the treatment response. We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, in comparison to a group with the same treatment responses but better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups. Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
Finally, we also compare our model against its unadjusted alternative (CNSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
Training.
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2. All models were optimised for 1000 epochs using an Adam optimiser [23] with early stopping.
Evaluation.
As, in the synthetic experiments, the subgroup structure is known, we measure the adjusted [^8] Random patient assignment results in an adjusted Rand-Index of 0. Rand-Index [39], which quantifies how the estimated assignment aligns with the known underlying group structure. Additionally, we use the integrated squared error (ISE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
[…]
where $\hat{\tau}_{k}$ is the estimated treatment effect for subgroup $k$, and $\tau_{k}$ is the ground truth. Our experiments use a numerical approximation of the integral at the last observed event time. Further, we measure the ISE at the population level, defined as
[…]
3.3 Treatment effect recovery
Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the "Observational" treatment assignment with the shaded area representing 95% CI. The log-likelihood presents an elbow around the underlying number of subgroups.
Recovering the underlying number of subgroups.
A limitation of our approach is that we need to choose the number of subgroups a priori. An important question to analyse is, therefore, whether we can identify the true number of subgroups with our model. Figure 4 presents the average NLL obtained by cross-validation for models with different numbers of subgroups $K$. The dotted lines represent the elbow heuristic [50], which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, aligning with the underlying generative process. This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which first model the treatment effect and subsequently cluster the population. Consequently, survival outcomes do not directly guide the choice of $K$ in classical approaches.
Discovering subgroups.
Next, we present quantitative results of our method against benchmarks. Table 1 presents the performance of the different methodologies on the two studied scenarios. Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival distributions.
[…]
Table 1: Cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation. ISEpop quantifies the error in estimating the average treatment effect, while ISEk reflects how the estimated treatment effects align with the ground truth. Best performance per column and simulation scenario are marked in bold. Our proposed CNSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
Our proposed CNSC methods outperform all alternatives of CMHE, the current state-of-the-art method, in the proposed experiments. We hypothesise CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log-hazards. Neither of these two assumptions is likely to hold in real-world settings. In contrast, our approach does not constrain the treatment effect through flexible modelling of the survival function under both treatment regimes.
Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower ISEpop) but still underperforms in comparison to our proposed methods. These experiments highlight the advantage of CNSC in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
Moreover, the Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. While all methodologies present larger ISE in the observational setting, we observe a larger performance gap between CMHE and CNSC, highlighting the importance of accounting for the treatment assignment bias. Further, the two CNSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. The superiority of CNSC over its unadjusted alternative is evident in the observational setting, where CNSC’s average treatment effect estimation performance is better, particularly in the recovery of treatment effect of subgroup 3, where not adjusting for treatment assignment doubles the ISE.
4 Related work
While survival subgrouping has been proposed through mixtures of distributions [32, 33, 19] to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current literature focuses on estimating population or individual treatment effects [9, 45, 22, 55]. While these models help to understand the average population-wide response to treatment and aim to estimate individualised treatment effects, they do not provide an understanding of the patient groups that may benefit more or be harmed more by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
Discovering intervenable subgroups is core to medical practice, particularly identifying subgroups of treatment effect, as patients do not respond like the average [6, 43, 44] and the average may conceal differential treatment responses. Identification of subgroups has long been used to design RCTs. Indeed, subgroups identified a priori can then be tested through trials [8, 41]. A posteriori analyses have gained traction to uncover subgroups of patients from existing RCTs to understand the underlying variability of responses.
The first set of subgroup analysis methodologies consists of a step-wise approach: (i) estimate the ITE and (ii) uncover subgroups using a second model to explain the heterogeneity in ITE. Foster et al. [11], Qi et al. [38] describe the virtual twins approach in which one models the outcome using a decision tree for each treatment group. The difference between these decision trees results in the estimated treatment effects. A final decision tree aims to explain these estimated treatment effects to uncover subgroups. Similar approaches have been explored with different meta-learners [53], or Bayesian additive models [17], or replacing the final step with a linear predictor to uncover the feature influencing heterogeneity [7]. However, Guelman et al. [12] discuss the drawbacks associated with these approaches. Notably, the two-step optimisation may not lead to recovery of the underlying subgroups of treatment effects.
Tree-based approaches have been proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. [49] introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey and Imbens [2, 1] improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager and Athey [51] agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect [27, 28]. Alternatively, McFowland III et al. [30] propose pattern detection and Wang and Rudin [52], causal rule set learning to uncover these subgroups. However, all these approaches rely on a criticised local optimisation criterion [26] and greedy split exploration. Recently, Nagpal et al. [31] addressed the local optimisation by constraining the treatment response shape to a linear form in a mixture of Cox models.
The previous approaches uncover subgroups of treatment effect but often consider RCTs with binary outcomes, not the observational setting with the inherently continuous survival outcomes that we are interested in. At the intersection with survival analysis, Zhang et al. [54] extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. [17] propose Bayesian additive models and Zhu and Gallego [56] propose a step-wise approach with propensity weighting to study observational data. Closest to our work, [21, 36] propose to uncover subgroups with survival outcomes. Specifically, Nagpal et al. [36] stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As discussed, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data. Similarly, Jia et al. [21] analyse RCTs with a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework.
5 Conclusion
This work introduces a novel methodology to uncover subgroups of treatment responses in observational time-to-event data. Our methodology addresses, to date, unaddressed challenges in subgroup discovery, specifically non-randomised treatment assignment and censoring. While modelling the observed outcome, the approach simultaneously identifies patient subgroups with different treatment responses in a joint approach. These groups may motivate the design of future clinical trials and support the development of guidelines to better target treatment to subgroups that would most benefit from them.
Our experiments demonstrate the capacity of our proposed method to uncover underlying groups of treatment effects. However, causal modelling relies on empirically unverifiable assumptions. The invalidity of these assumptions is a lesser concern in hypothesis generation, as offered by the proposed model, than in individualised estimates aimed at treatment recommendations.
Serving the purpose of a hypothesis-generating tool, we invite practitioners to further study the subgroups, identified in observational studies, in RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
Acknowledgments and Disclosure of Funding
VJ acknowledges the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2) and the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1. FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z), and the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
Appendix A Proof
A.1 Individualised Treatment Effect
This section derives the individualised treatment effect expression introduced in Eq. (3).
$\tau(t,x):$ $=\mathbb{E}(\mathbbm{1}(T_{1}\geq t)-\mathbbm{1}(T_{0}\geq t)\mid X=x)$
$=\mathbb{E}(\mathbbm{1}(T_{1}\geq t)\mid X=x)-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
$=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under Assumption 4 and 5)
A.2 Upper-Bound Simplification
In Section 2, we claim that under the considered DAG presented in Figure 2,
$\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$
with $A$, being the treatment assignment. The following derives this result:
$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$ (Under Assumption 3)
with $q$, the density function. From this expression, $q_{Z\mid X}^{A=0}=q_{Z\mid X}^{A=1}$, which results in the distance between these distributions being null.
Appendix B Synthetic analysis
B.1 Data generation
We consider a synthetic population of $N=30,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups. The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to [34]. The following describes our generation process:
Covariates.
Each patient’s membership $Z$ is drawn from a multinomial with equal probability. Group membership informs the two first covariates through the parametrisation of the bivariate normal distribution with centres $c_{k}$ equal to $(0,2.25)$, $(-2.25,-1)$, and $(2.25,-1)$. All other covariates are drawn from standard normal distributions. Formally, this procedure is described as:
$Z$ $\sim\text{Mult}\left(1,\left[\frac{1}{3},\frac{1}{3},\frac{1}{3}\right]\right)$
$X_{[1,2]}\mid Z=k$ $\sim\text{MVN}(c_{k},I^{2})$
$X_{[3:10]}$ $\sim\text{MVN}(0,I^{8})$
with MVN denoting a multivariate normal distribution, and $I^{n}$, the identity covariance matrix of dimension $n$.
Treatment response.
For each subgroup, event times under treatment and control regimes are drawn from Gompertz distributions, with parameters that are functions of group-specific coefficients ($B^{0}$ and $\Gamma^{0}$ for the event time when untreated and $B^{1}$ and $\Gamma^{1}$ when treated) and the patient’s covariates.
$B^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$\Gamma^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$T_{0}\mid Z,X,B^{0}_{z},\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
$B^{1}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$\Gamma^{1}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{1}_{z}=(z,x,\beta_{z}^{1},\gamma_{z}^{1})$ $\sim\text{Gompertz}\left(w_{1}(\beta_{z}^{1},x),s_{1}(\gamma_{z}^{1},x)\right)$
with $w_{0}$, $w_{1}$ two functions parametrising the Gompertz distributions’ shape as $w_{0}(\beta,x):=|\beta[0]|+(x[5:10]\cdot\beta[5:10])^{2}$, $w_{1}(\beta,x):=|\beta[0]|+(x[1:5]\cdot\beta[1:5])^{2}$, and the shift parameter parameterised as $s_{0}(\gamma,x):=|\gamma[0]|+|(x[1:5]\cdot\gamma[1:5])|$ and $s_{1}(\gamma,x):=|\gamma[0]|+|(x[5:10]\cdot\gamma[5:10])|$ where $v[i]$ described the $i^{th}$ element of the vector $v$. These functions aim to introduce non-linear responses with discrepancies between control and treatment regimes. Note that we allow covariates to influence the survival distribution as a patient’s covariates influence Gompertz’s shapes and scales.
Treatment assignment.
The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world applications. After drawing the treatment assignment probability $P$, we assign each patient to a given treatment. We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as "Randomised" and "Observational". "Randomised" consists of a Bernoulli draw using the realisation of $P$. "Observational" reflects an assignment dependent upon the observed covariates.
$P$ $\sim\text{Uniform}(0.25,0.75)$
$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(p)$
$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times p)$
with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$. In our experiment, we chose $\Phi(x)=x[1]^{2}+x[2]^{2}$.
Censoring.
Finally, our work focuses on right-censored data. To generate censoring independent of the treatment and event, we draw censoring from another Gompertz distribution as follows:
$B^{C}$ $\sim\text{MVN}(0,I^{5})$
$C\mid X,B^{C}=(x,\beta)$ $\sim\text{Gompertz}\left(w_{c}(\beta,x),0\right)$
$T^{\prime}$ $=A\cdot T_{1}+(1-A)\cdot T_{0}$
$T$ $=\min(C,T^{\prime})$
$D$ $=\mathbbm{1}(C>T^{\prime})$
with $w_{c}:=(x[5:10]\cdot\beta)^{2}$, the scale of the censoring Gompertz distribution.
From these data, the goal is to model the treatment effect from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)*t_{0}+a*t_{1})$ and $D=\mathbbm{1}_{C>T}$.
B.2 Optimisation grid search
We adopted a 10-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[\![25,50,100]\!]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $250$ or $500$.
Appendix C Case Study: SEER dataset
To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^9] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017. Following [25, 10, 20], we select women who died from the condition or from cardiovascular diseases. From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes. To this end, we subselect patients with recorded treatment and who received chemotherapy. These criteria led to the selection of 239,855 patients with 22 tumour characteristics measured at diagnosis.
Selection of $K$.
From the selected population, our aim is to identify heterogeneous responses to adjuvant treatment. The first challenge is the selection of the number of groups to use ($K$). We advise to follow medical actionability and consider the change in treatment effects and size of the subgroups when increasing this parameter. In the absence of experts’ intuition, one may rely on an elbow rule heuristic over $l_{F}$. This means selecting the number of subgroups after which one observes limited likelihood improvement as used in [19]. Using Figure 5, the negative log-likelihood presents an elbow for $K=2$.
Figure 5: Cross-validated negative log-likelihood as a function of the number of groups ($K$).
Treatment effect subgroups.
In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy. This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist [24], making this setting more likely to meet the positivity assumption (Assumption 6), necessary to study causality in observational data.
(a) Neural Survival Treatment
(b) Cox Mixture of Heterogenous Effects
Figure 6: Averaged treatment effect subgroups across 5-fold cross-validation observed in the Seer dataset with the shaded areas representing 95% CI.
Uncovering treatment response.
Figure 6 presents the identified treatment effect subgroups when using CNSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning. As shown by Figure 5, using the previously described elbow rule leads to the same number of subgroups. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
Using a permutation test, we identify as the covariates that most impact the likelihood associated with the model. Figure 7 displays the 10 covariates most indicative of the different treatment response subgroups.
Figure 7: Causal Neural Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
Table 2 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) [42]. Both methodologies identify a population with limited treatment response. However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.00 (0.00) 93.3% 55.5% 1.18 (5.74) 17.4% 46.6% Subgroup 1 0.62 (0.16) 6.7% 47.5% 17.75 (16.08) 23.4% 48.6%
Table 2: Causal Neural Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy. However, our methodologies remain hypothesis-generating tools, requiring further experimental validation, particularly due to potential confounding through hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options.

### v3

@@
 Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials.
-Existing approaches for subgroup analysis primarily focus on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised.
-Furthermore, the patient cohort of an RCT is often constrained by cost, and is not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice.
-Therefore, when applied to observational studies, such approaches suffer from significant statistical biases because of the non-randomisation of treatment.
+Existing approaches for subgroup analysis primarily rely on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised.
+RCTs’ patient cohorts are often constrained by cost, rendering them not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice.
+When applied to observational studies, subgroup analysis approaches suffer from significant statistical biases particularly because of the non-randomisation of treatment.
 Our work introduces a novel, outcome-guided method for identifying treatment response subgroups in observational studies.
@@
 In experiments, our approach significantly outperforms the current state-of-the-art method for outcome-guided subgroup analysis in both randomised and observational treatment regimes.
-Data and Code Availability
-Experiments are performed on a synthetic dataset and a real-world dataset from the Surveillance, Epidemiology, and End Results program [^1] Available at https://seer.cancer.gov/.
-For reproducibility, implementation of the method, synthetic dataset, and analyses are available on Github  [^2] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
 1 Introduction
 Understanding heterogeneous therapeutic responses between patient subgroups is the core of treatment guidelines and the development of new drugs.
-Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed [11].
-For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes [18], shaping subsequent guidelines.
+Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed (Foster et al., 2011).
+For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes (Investigators, 1996), shaping subsequent guidelines.
 Figure 1 illustrates this idea: two groups may present opposite treatment responses, and would benefit from different recommendations.
@@
 In RCTs, patients are randomly assigned to control or treated groups, allowing researchers to assess the impact of an intervention or treatment.
-However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies [16].
-Our work introduces a novel approach that diverges from traditional RCT-based methodologies [34, 36] by leveraging routinely collected observational data to identify patient subgroups.
+However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies (Hernán & Robins, 2016).
+Our work introduces a novel approach that diverges from traditional RCT-based methodologies (Nagpal et al., 2022a; 2023) by leveraging routinely collected observational data to identify patient subgroups.
 Observational studies encompass larger and more diverse cohorts reflective of real-world practices, offering the potential to uncover subgroups of treatment responses, that could be missed in RCTs.
-Prior works [5, 9, 29] have leveraged observational data while addressing biases from non-random treatment assignments [4, 15, 14].
+Prior works (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017) have leveraged observational data while addressing biases from non-random treatment assignments (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018).
 However, this body of literature primarily focuses on estimating (i) averaged treatment effects at the population level or (ii) individualised treatment effects, thereby overlooking the identification of treatment effect subgroups.
@@
 Contributions.
-Our work extends Neural Survival Clustering [19] to estimate treatment effects, introducing the novel Causal Neural Survival Clustering (CNSC).
-This methodology is the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects.
+Our work introduces the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects.
 Section 2 formalises the problem of treatment subgroup discovery and non-randomisation correction.
 Then, we introduce the proposed monotonic neural network implementation in Section 2.
-Finally, we evaluate the proposed methodology on two datasets in Section 3 and compare it with related approaches described in Section 4.
+Finally, we evaluate the proposed methodology on two synthetic datasets in Section 3, an extensive set of alternative settings in App.
+B and a real-world example described in App.
+C and compare it with related approaches reviewed in Section 4.
 2 Method
-In this section, we describe our method, referring to Appendix A for technical proofs.
 2.1 Problem setup
@@
 Following Jeanselme et al.
-[19], we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^3] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3..
+(2022), we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3..
 We assume that $Z$ depends on $X$ and influences $T^{\prime}$.
@@
 Likelihood-based model fitting.
-One can maximise the log-likelihood associated with the observed realisation $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i$ (we omit $i$ where redundant) to estimate the survival function (Eq.
-(1)).
-Under Assumption 1, Equation (2) describes this factual [^4] Referring to the likelihood of the observed (rather than counterfactual) realisations.
+One can maximise the log-likelihood associated with the observed realisations $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i\in[\![1,N]\!]$ with $N$ the total number of patients (we omit $i$ where redundant) to estimate the survival function (Eq.
+equation 1).
+Under Assumption 1, Equation equation 2 describes this factual [^2] Referring to the likelihood of the observed (rather than counterfactual) realisations.
 log-likelihood.
@@
 where $\theta$ is the set of parameters characterising the estimated survival function $S_{\theta}$.
-Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see [19] for derivation).
+Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see Jeanselme et al.
+(2022) for derivation).
 Each censored patient ($d_{i}=0$) contributes the probability of not observing an event before $t_{i}$, i.e., $S_{\theta}(t_{i}\mid x_{i})$, under assumption 1 of independence between the censoring time and the event of interest conditional on the covariates.
@@
 The treatment assignment $A$ is independent of the subgroup membership $Z$ given the covariates $X$. Formally, $A\perp\!\!\!\perp Z\mid X$.
-Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$).
-Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
-$Z$ stands for the latent subgroup structure explaining the event time under no treatment $T^{\prime}_{0}$ and under treatment $T^{\prime}_{1}$.
-$A$ is the assigned treatment.
-$C$ is the censoring time.
-This assumption is realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) subgroups are unknown and would consequently not inform treatment.
-If subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
+This assumption relaxes the traditional assumption of treatment randomisation.
+Note that this assumption remains realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) if subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
+Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
 Further, the following three assumptions are necessary to estimate treatment effects and are common in the causal literature.
@@
 For example, clinicians might recommend more aggressive treatment to patients in more severe conditions.
-This absence of randomisation results in a covariate shift between the treated and non-treated populations [9] as their covariate distributions differ [6].
+This absence of randomisation results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021).
 In this setting, estimating survival by maximising the factual likelihood is no-longer enough to estimate the counterfactual survival distribution.
-Under Assumption 6, approaches such as re-weighting [48], or penalisation on the dissimilarity of learnt representations [22, 47], or a combination of these approaches [9, 13, 45] are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest.
+Under assumption 6, approaches such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity of learnt representations (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Shalit et al., 2017) are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest.
 Specifically, Shalit et al.
-[45] demonstrate that the negative log likelihood is upper-bounded by observable quantities.
+(2017) demonstrate that the negative log likelihood is upper-bounded by observable quantities.
 Here we extend their notations to the survival setting as
 $-l(\theta)\leq-l_{F}^{*}(\theta)+\gamma\cdot\text{IPM}(q_{\Phi}^{A=0},q_{\Phi}^{A=1}),$ (4)
-where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in Eq.(2)) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
+where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in equation 2) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
 Shalit et al.
@@
 A.2 for derivation).
-As a consequence, our work focuses on the first term of the upper bound in Eq.
-(4) by re-weighting the factual likelihood with the patient’s propensity score [13, 45].
-Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting [3] scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
+This assumption is consequently necessary to correctly estimate the subgroups’ treatment effect, without penalising the clustering structure.
+As a consequence, our work focuses on the first term of the upper bound in equation 4 by re-weighting the factual likelihood with the patient’s propensity score (Hassanpour & Greiner, 2019; Shalit et al., 2017).
+Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting (Austin & Stuart, 2015) scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
 $w_{i}^{-1}=\begin{cases}0.05&\text{ if }\hat{p}_{A}(x_{i})<0.05\\ 0.95&\text{ if }\hat{p}_{A}(x_{i})>0.95\\ a_{i}\cdot\hat{p}_{A}(x_{i})+(1-a_{i})\cdot(1-\hat{p}_{A}(x_{i}))&\text{ otherwise},\end{cases}$ (5)
 where $i$ is the patient index, and $a_{i}$ is the realisation of $A$ for patient $i$.
-Using the factual likelihood from Eq.
-(2) and the weights $w_{i}$ defined in Eq.
-(5), we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
+Using the factual likelihood from equation 2 and the weights $w_{i}$ defined in equation 5, we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
 $\sum_{i,d_{i}=1}w_{i}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i},a_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}w_{i}\log S_{\theta}(t_{i}\mid x_{i},a_{i})$ (6)
-Figure 3: Causal Neural Survival Clustering architecture.
+Figure 3: Causal Survival Clustering architecture.
 Latent parameter $l_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the cumulative hazard $\Lambda_{k}$ under both treatment regimes.
@@
 Subgroup assignment.
-Similar to [19], a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
+Similar to Jeanselme et al.
+(2022), a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
 $G(x):=[\mathbb{P}(Z=k\mid x)]_{k=1}^{K}$
@@
 Each subgroup $k$ is represented by a vector $l_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the cumulative hazard functions.
-The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^5] To ensure this constraint, we apply a square function on all weights as proposed in [19, 20]., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$.
-We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks [46]:
+The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in Jeanselme et al.
+(2022; 2023)., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$.
+We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
 $\Lambda_{k}(t):=t\cdot M(l_{k},t)$
-By modelling these two cumulative intensity functions, we can estimate the Group Average Treatment Effect (GATE), $\hat{\tau}_{k}(t)$, in subgroup $k$ as
+By modelling these two cumulative intensity functions, we can estimate the Subgroup Average Treatment Effect, $\hat{\tau}_{k}(t)$, in subgroup $k$ as
 $\hat{\tau}_{k}(t):=\mathbb{E}[\tau(t,X)\mid Z=k)]=\exp(-\Lambda_{k}(t\mid A=1))-\exp(-\Lambda_{k}(t\mid A=0))$
@@
 Under treatment randomisation, as in RCTs, the previous components would accurately model the observed survival outcomes and identify subgroups of treatment effects by maximising the factual likelihood.
-As previously discussed, to account for the treatment non-randomisation in observational studies, we weigh the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
+As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
 $W(x):=\mathbb{P}(A=1\mid x)$
@@
 First, the network $W$ is trained to predict the binary treatment assignment by minimising the cross-entropy of receiving treatment.
-Then, training all other components relies on minimising the weighted factual log-likelihood introduced in Eq.
-(6).
-The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative used in the log likelihood [19, 20, 40].
+Then, training all other components relies on minimising the weighted factual log-likelihood introduced in equation 6.
+The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes readily provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative in the log likelihood (Jeanselme et al., 2022; 2023; Rindt et al., 2022).
 3 Experimental analysis
 As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known.
-Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer.
-Our code for the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^6] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
+Appendix C accompanies these synthetic results with the analysis of heterogeneity in real-world, adjuvant radiotherapy responses for patients diagnosed with breast cancer.
+Our code to produce the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
 3.1 Data generation
-We generate a population of $30,000$ divided into $K=3$ subgroups.
-We draw 10 covariates from normal distributions and survival distributions for each treatment regime following Gompertz distributions [37] parameterised by group membership and individual covariates.
+We generate a population of $3,000$ equally divided into $K=3$ subgroups.
+We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parameterised by group membership and individual covariates.
+Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model.
 This choice reflects a setting in which time-to-event distributions under treatment or control regimes are different functions of the covariates.
-This simulation is more likely to capture the complexity of real-world responses, contrary to traditional evaluations of subgroup analysis, which often assume a linear treatment response, i.e., a simple shift between treated and untreated distributions.
-Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as observed in observational studies (Observational).
-Finally, non-informative censoring times are drawn also following a Gompertz distribution.
+This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis which often assume a linear treatment response, i.e., a shift between treated and untreated distributions.
+Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational).
+Finally, non-informative censoring times are drawn following a Gompertz distribution.
 Details of the generative process of our synthetic dataset are deferred to App.
@@
 3.2 Empirical settings
-Benchmarks.
-We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE [34] [^7] Implemented in the Auton-Survival library [35]) which uncovers treatment effect and baseline survival latent groups.
+Benchmark methods.
+We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE Nagpal et al.
+(2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups.
 This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted.
@@
 Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
-Finally, we also compare our model against its unadjusted alternative (CNSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
-Training.
-We perform a 5-fold cross-validation for both Randomised and Observational simulations.
-For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2.
-All models were optimised for 1000 epochs using an Adam optimiser [23] with early stopping.
+Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
+Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately.
+First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in Nagpal et al.
+(2022a), referred as Kmeans + TE in the following.
+Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters.
+However, this approach is limited by its implicit assumption of randomisation and its pre-specified number of subgroups, which is not informed by observed outcomes.
+Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them.
+As proposed in the literature, we propose to use a survival tree to estimate response under each treatment regime and then use a KMeans on the estimated treatment effects, computed as the difference between survival estimates.
+Note that this methodology also assumes treatment randomisation, and the clustering is uninformed by the covariates.
+For details on training and hyperparameter optimisation, we refer to App.
+B.2.
 Evaluation.
-As, in the synthetic experiments, the subgroup structure is known, we measure the adjusted [^8] Random patient assignment results in an adjusted Rand-Index of 0.
-Rand-Index [39], which quantifies how the estimated assignment aligns with the known underlying group structure.
-Additionally, we use the integrated squared error (ISE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
+In the synthetic experiments, the subgroup structure is known.
+We measure the adjusted [^6] Random patient assignment results in an adjusted Rand-Index of 0.
+Rand-Index (Rand, 1971), which quantifies how the estimated assignment aligns with the known underlying group structure.
+Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
 […]
-where $\hat{\tau}_{k}$ is the estimated treatment effect for subgroup $k$, and $\tau_{k}$ is the ground truth.
-Our experiments use a numerical approximation of the integral at the last observed event time.
-Further, we measure the ISE at the population level, defined as
-[…]
+where $\hat{\tau}_{\hat{k}}$ is the estimated treatment effect for subgroup $\hat{k}=\operatorname*{arg\,max}_{l}\mathbb{E}_{x\in k}(Z=l\mid x)$, i.e.
+the most likely assigned cluster for patients in the underlying $k$ cluster, and $\tau_{k}$ is the ground truth.
 3.3 Treatment effect recovery
+In this section, we present the results under the previously described data generation process. App. B.4 presents alternative datasets which demonstrate the flexibility of the proposed methodology.
 Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the "Observational" treatment assignment with the shaded area representing 95% CI.
@@
 Recovering the underlying number of subgroups.
-A limitation of our approach is that we need to choose the number of subgroups a priori.
-An important question to analyse is, therefore, whether we can identify the true number of subgroups with our model.
+For all methodologies, we need to choose the number of subgroups a priori.
+An important question is, therefore, whether we can identify the true number of subgroups with our model.
 Figure 4 presents the average NLL obtained by cross-validation for models with different numbers of subgroups $K$.
-The dotted lines represent the elbow heuristic [50], which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood.
-Using this heuristic, the optimal choice for $K$ is $3$, aligning with the underlying generative process.
-This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which first model the treatment effect and subsequently cluster the population.
-Consequently, survival outcomes do not directly guide the choice of $K$ in classical approaches.
+The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood.
+Using this heuristic, the optimal choice for $K$ is $3$, which agrees with the underlying generative process.
+App.
+B.3 explores the sensitivity to this choice, demonstrating the method’s robustness.
+This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses which separate clustering from treatment effect estimates.
+In these methods, survival outcomes cannot directly guide the choice of $K$.
 Discovering subgroups.
-Next, we present quantitative results of our method against benchmarks.
-Table 1 presents the performance of the different methodologies on the two studied scenarios.
+Next, we present quantitative results of our method against the benchmarks methods.
+Table 1 presents the performance of the different methodologies under the two studied scenarios.
 Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival distributions.
-[…]
-Table 1: Cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation.
-ISEpop quantifies the error in estimating the average treatment effect, while ISEk reflects how the estimated treatment effects align with the ground truth.
-Best performance per column and simulation scenario are marked in bold.
-Our proposed CNSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
-Our proposed CNSC methods outperform all alternatives of CMHE, the current state-of-the-art method, in the proposed experiments.
-We hypothesise CMHE’s parametrisation and assumptions explain this difference.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.798 (0.054) 0.042 (0.012) 0.032 (0.007) 0.015 (0.006) CSC Unadjusted 0.804 (0.052) 0.032 (0.009) 0.027 (0.007) 0.018 (0.010) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.258 (0.155) - 0.088 (0.007) - CMHE ($K=L$) 0.392 (0.054) 0.338 (0.262) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.812 (0.038) 0.037 (0.013) 0.034 (0.009) 0.022 (0.005) CSC Unadjusted 0.727 (0.084) 0.048 (0.019) 0.025 (0.011) 0.025 (0.009) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
+Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation.
+Best performance per column and simulation scenario are marked in bold, second best in italic.
+’-’ describes when the methodology diverges.
+Our proposed CSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
+CSC outperforms all CMHE alternative, the current state-of-the-art method.
+CMHE’s parametrisation and assumptions explain this difference.
 Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log-hazards.
-Neither of these two assumptions is likely to hold in real-world settings.
-In contrast, our approach does not constrain the treatment effect through flexible modelling of the survival function under both treatment regimes.
-Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower ISEpop) but still underperforms in comparison to our proposed methods.
-These experiments highlight the advantage of CNSC in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
-Moreover, the Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation.
-While all methodologies present larger ISE in the observational setting, we observe a larger performance gap between CMHE and CNSC, highlighting the importance of accounting for the treatment assignment bias.
-Further, the two CNSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context.
-The superiority of CNSC over its unadjusted alternative is evident in the observational setting, where CNSC’s average treatment effect estimation performance is better, particularly in the recovery of treatment effect of subgroup 3, where not adjusting for treatment assignment doubles the ISE.
+Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data.
+In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes.
+Increasing the number of subgroups, as shown in $K=L$ improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods.
+These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
+CSC presents the best performance in identifying the underlying subgroup.
+Our proposed CSC method outperforms all approaches on the Rand-Index.
+In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are aligned.
+KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index.
+Consequently, the identified clusters differ in their treatment responses as shown by the large IAE.
+The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE.
+However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates.
+This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
+Treatment assignment correction improves subgroup identification in observational settings.
+The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation.
+The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context.
+Note that estimating propensity slightly reduces performance due to error introduced, reducing performance for the corrected alternative.
+In a randomised setting, CSC Unadjusted should be preferred.
+However, the superiority of CSC over its unadjusted alternative is evident in the observational setting, where CSC better recovers the different groups as shown by the Rand-Index and subgroup treatment effects.
+Finally, Appendices B.4.1 to B.4.5 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
+Real-world heterogeneity analysis in adjuvant radiotherapy responses (App. C).
+Using the Surveillance, Epidemiology, and End Results program [^7] Available at https://seer.cancer.gov/(Seer), App.
+C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research.
+Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment.
+Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis.
+Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
 4 Related work
-While survival subgrouping has been proposed through mixtures of distributions [32, 33, 19] to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse.
-The current literature focuses on estimating population or individual treatment effects [9, 45, 22, 55].
+While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a; b; Jeanselme et al., 2022) to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse.
+The current literature focuses on estimating population or individual treatment effects (Curth et al., 2021; Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020).
 While these models help to understand the average population-wide response to treatment and aim to estimate individualised treatment effects, they do not provide an understanding of the patient groups that may benefit more or be harmed more by treatment.
 Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
-Discovering intervenable subgroups is core to medical practice, particularly identifying subgroups of treatment effect, as patients do not respond like the average [6, 43, 44] and the average may conceal differential treatment responses.
+Discovering intervenable subgroups is core to medical practice, particularly identifying subgroups of treatment effect, as patients do not respond like the average (Bica et al., 2021; Ruberg et al., 2010; Sanchez et al., 2022) and the average may conceal differential treatment responses.
 Identification of subgroups has long been used to design RCTs.
-Indeed, subgroups identified a priori can then be tested through trials [8, 41].
+Indeed, subgroups identified a priori can then be tested through trials (Cook et al., 2004; Rothwell, 2005).
 A posteriori analyses have gained traction to uncover subgroups of patients from existing RCTs to understand the underlying variability of responses.
@@
 Foster et al.
-[11], Qi et al.
-[38] describe the virtual twins approach in which one models the outcome using a decision tree for each treatment group.
+(2011); Qi et al.
+(2021) describe the virtual twins approach in which one models the outcome using a decision tree for each treatment group.
 The difference between these decision trees results in the estimated treatment effects.
 A final decision tree aims to explain these estimated treatment effects to uncover subgroups.
-Similar approaches have been explored with different meta-learners [53], or Bayesian additive models [17], or replacing the final step with a linear predictor to uncover the feature influencing heterogeneity [7].
+Similar approaches have been explored with different meta-learners (Xu et al., 2023), or Bayesian additive models (Hu et al., 2021), or replacing the final step with a linear predictor to uncover the feature influencing heterogeneity (Chernozhukov et al., 2018).
 However, Guelman et al.
-[12] discuss the drawbacks associated with these approaches.
+(2015) discuss the drawbacks associated with these approaches.
 Notably, the two-step optimisation may not lead to recovery of the underlying subgroups of treatment effects.
-Tree-based approaches have been proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect.
+Tree-based approaches were proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect.
 Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects.
 Su et al.
-[49] introduce a recursive population splitting based on the average difference in treatment effect between splits.
-Athey and Imbens [2, 1] improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation.
-Wager and Athey [51] agglomerate these causal trees into causal forests for improved ITE estimation.
-Each obtained split in the decision tree delineates two subgroups of treatment effect [27, 28].
+(2009) introduce a recursive population splitting based on the average difference in treatment effect between splits.
+Athey & Imbens (2015; 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation.
+Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation.
+Each obtained split in the decision tree delineates two subgroups of treatment effect Lipkovich et al.
+(2011); Loh et al.
+(2015).
 Alternatively, McFowland III et al.
-[30] propose pattern detection and Wang and Rudin [52], causal rule set learning to uncover these subgroups.
-However, all these approaches rely on a criticised local optimisation criterion [26] and greedy split exploration.
+(2018) propose pattern detection and Wang & Rudin (2022), causal rule set learning to uncover these subgroups.
+However, all these approaches rely on a local optimisation criterion (Lipkovich & Dmitrienko, 2014) and greedy split exploration.
 Recently, Nagpal et al.
-[31] addressed the local optimisation by constraining the treatment response shape to a linear form in a mixture of Cox models.
-The previous approaches uncover subgroups of treatment effect but often consider RCTs with binary outcomes, not the observational setting with the inherently continuous survival outcomes that we are interested in.
+(2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
+Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with inherently continuous survival outcomes that our paper explores.
 At the intersection with survival analysis, Zhang et al.
-[54] extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves.
+(2017) extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves.
 Similarly, Hu et al.
-[17] propose Bayesian additive models and Zhu and Gallego [56] propose a step-wise approach with propensity weighting to study observational data.
-Closest to our work, [21, 36] propose to uncover subgroups with survival outcomes.
-Specifically, Nagpal et al.
-[36] stratify the population into three groups: non-, positive- and negative responders to treatment.
+(2021) propose Bayesian additive models and Zhu & Gallego (2020) propose a step-wise approach with propensity weighting to study observational data.
+Closest to our work, (Jia et al., 2021; Nagpal et al., 2023) propose to uncover subgroups within RCTs with survival outcomes.
+Jia et al.
+(2021) propose a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework.
+Similarly, Nagpal et al.
+(2023) stratify the population into three groups: non-, positive- and negative responders to treatment.
 An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect.
-As discussed, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
-Similarly, Jia et al.
-[21] analyse RCTs with a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework.
+As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
 5 Conclusion
-This work introduces a novel methodology to uncover subgroups of treatment responses in observational time-to-event data.
-Our methodology addresses, to date, unaddressed challenges in subgroup discovery, specifically non-randomised treatment assignment and censoring.
-While modelling the observed outcome, the approach simultaneously identifies patient subgroups with different treatment responses in a joint approach.
-These groups may motivate the design of future clinical trials and support the development of guidelines to better target treatment to subgroups that would most benefit from them.
-Our experiments demonstrate the capacity of our proposed method to uncover underlying groups of treatment effects.
+This work fills the current gap in the literature of methodologies to identify subgroups of treatment responses in observational time-to-event data.
+While modelling the observed outcome, the approach jointly identifies patient subgroups with different treatment responses.
+Our experiments demonstrate the capacity of our proposed method to uncover groups of treatment effects in both synthetic and real-world datasets.
 However, causal modelling relies on empirically unverifiable assumptions.
-The invalidity of these assumptions is a lesser concern in hypothesis generation, as offered by the proposed model, than in individualised estimates aimed at treatment recommendations.
-Serving the purpose of a hypothesis-generating tool, we invite practitioners to further study the subgroups, identified in observational studies, in RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
-Acknowledgments and Disclosure of Funding
-VJ acknowledges the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2) and the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
-FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z), and the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
+The invalidity of these assumptions is a lesser concern in our targeted setting of hypothesis generation, in comparison to individual estimation of treatment recommendations.
+Serving the purpose of a hypothesis-generating tool, we invite practitioners to further investigate the subgroups identified in observational studies through RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
+Acknowledgements
+The authors acknowledge the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2).
+VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
+FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z).
 Appendix A Proof
 A.1 Individualised Treatment Effect
-This section derives the individualised treatment effect expression introduced in Eq. (3).
+This section derives the individualised treatment effect expression introduced in equation 3.
 $\tau(t,x):$ $=\mathbb{E}(\mathbbm{1}(T_{1}\geq t)-\mathbbm{1}(T_{0}\geq t)\mid X=x)$
@@
 $\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$
-with $A$, being the treatment assignment. The following derives this result:
-$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$ (Under Assumption 3)
+with $A$, being the treatment assignment. Under assumption 3, expressing this quality corresponds to:
+$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$
 with $q$, the density function. From this expression, $q_{Z\mid X}^{A=0}=q_{Z\mid X}^{A=1}$, which results in the distance between these distributions being null.
@@
 B.1 Data generation
-We consider a synthetic population of $N=30,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups.
-The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to [34].
+We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups.
+The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to Nagpal et al.
+(2022a).
 The following describes our generation process:
@@
 $X_{[1,2]}\mid Z=k$ $\sim\text{MVN}(c_{k},I^{2})$
-$X_{[3:10]}$ $\sim\text{MVN}(0,I^{8})$
-with MVN denoting a multivariate normal distribution, and $I^{n}$, the identity covariance matrix of dimension $n$.
+$Y$ $\sim\text{Mult}\left(1,\left[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}\right]\right)$
+$X_{[3:10]}\mid Y=k$ $\sim\text{MVN}(c^{\prime}_{k},I^{8})$
+with MVN denoting a multivariate normal distribution, $c^{\prime}_{k}$ random cluster centres, and $I^{n}$, the identity covariance matrix of dimension $n$.
+Note that we introduce $Y$ to present a covariates structure that is independent of the treatment responses.
 Treatment response.
@@
 The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world applications.
-After drawing the treatment assignment probability $P$, we assign each patient to a given treatment.
+Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment.
 We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as "Randomised" and "Observational".
@@
 "Observational" reflects an assignment dependent upon the observed covariates.
-$P$ $\sim\text{Uniform}(0.25,0.75)$
-$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(p)$
-$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times p)$
+$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(0.5)$
+$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times 0.5)$
 with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$.
-In our experiment, we chose $\Phi(x)=x[1]^{2}+x[2]^{2}$.
+In our experiment, we chose $\Phi(x)=\sum x^{2}$.
 Censoring.
@@
 From these data, the goal is to model the treatment effect from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)*t_{0}+a*t_{1})$ and $D=\mathbbm{1}_{C>T}$.
-B.2 Optimisation grid search
-We adopted a 10-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[\![25,50,100]\!]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $250$ or $500$.
-Appendix C Case Study: SEER dataset
+B.2 Training and hyperparameter optimisation
+Training.
+We perform a 5-fold cross-validation for both Randomised and Observational simulations.
+For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2.
+All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
+Hyperparameter optimisation.
+We adopted a 100-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[10,25,50]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $100$ or $250$.
+B.3 Sensitivity
+In Section 3, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$.
+This section explores how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters.
+This analysis examines the risk associated with misspecifying the number of clusters.
+Figure 5 illustrates the cross-validated subgroup treatment effects when trained with $K=2,3$ and $4$ clusters.
+Uncertainty increases with misspecified models due to the instability of the identified subgroups.
+(a) $K=2$
+(b) $K=3$ (as in the main text)
+(c) $K=4$
+Figure 5: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross validated average estimated treatment effect. Lines in grey corresponds to the ground truth.
+B.4 Alternative data generations
+In this section, we explore alternative data generations to demonstrate the robustness of the proposed strategy.
+B.4.1 Unequal cluster size
+Section B.1 presents a population equally distributed over the different clusters.
+In medical settings, the underlying subgroups may differ in size.
+This section explores an alternative scenario in which the population is distributed over the three clusters: 62.5%, 25% and 12.5%.
+Table 2 summarises the performance in this simulation, echoing the main text’s conclusions.
+The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects..
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.637 (0.068) 0.064 (0.013) 0.021 (0.009) 0.062 (0.010) CSC Unadjusted 0.651 (0.049) 0.055 (0.009) 0.017 (0.010) 0.067 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.780 (0.086) 0.064 (0.027) 0.040 (0.011) 0.069 (0.010) CSC Unadjusted 0.616 (0.146) 0.072 (0.020) 0.039 (0.028) 0.117 (0.033) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
+Table 2: Cross-validated performance (with standard deviation in parenthesis) when clusters are of different sizes.
+B.4.2 Number of points
+Section B.1 describes a data generation with $3,000$ patients.
+This section presents results when there are $300$ and $30,000$ patients.
+A smaller number of patients may result in a larger impact of treatment non-randomisation but also impact the neural network capacity to identify underlying subgroups of treatment effect.
+A larger number makes non-randomisation less of a concern.
+Table 3 presents the performance with these different population sizes under an observational treatment assignment.
+A first observation is that all methodologies present better performance with larger number of points.
+Further, baselines that do not account for the assignment mechanisms present lower performance with $N=300$ as non-randomisation has an increased impact on treatment effect estimate with a smaller population.
+This difference decreases when $N=30,000$ with the virtual twins approach presenting the best recovery of the treatment effects.
+However, throughout the different settings, the proposed CSC presents the best Rand-Index indicating a good recovery of the underlying structure.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.504 (0.139) 0.131 (0.090) 0.085 (0.033) 0.091 (0.062) CSC Unadjusted 0.449 (0.106) 0.202 (0.049) 0.162 (0.055) 0.125 (0.088) CMHE ($K=3$) 0.126 (0.084) 0.340 (0.050) 0.135 (0.043) 0.166 (0.050) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.087 (0.118) 0.284 (0.056) 0.639 (0.939) 0.747 (1.014) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.774 (0.015) 0.018 (0.010) 0.012 (0.003) 0.030 (0.009) CSC Unadjusted 0.737 (0.022) 0.033 (0.006) 0.010 (0.002) 0.039 (0.008) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
+Table 3: Cross-validated performance (with standard deviation in parenthesis) with varying $N$ under observational treatment setting.
+B.4.3 Number of clusters
+Section B.1 describes a data generation with $K=3$ clusters.
+This section explores the behaviour of the different methodologies with an increased number of clusters $K=5$.
+Following the same data generation, we introduces two additional clusters with centres [-3, 3] and [4, 4].
+Table 4 presents the different models’ performance in this setting, echoing the same results as the main text.
+The proposed method best recovers the underlying clustering structure as shown by the Rand-Index.
+CMHE suffers from its inherent association between treatment and survival.
+The step-wise approach is unable to identify the clusters of interest if the data presents a structure that is not fully aligned with the outcome of interest.
+This scenario further validates our findings, demonstrating the methodology’s capacity to generalise beyond the scenario presented in the main text.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ Randomised CSC 0.475 (0.038) 0.073 (0.030) 0.059 (0.023) 0.041 (0.015) 0.099 (0.024) 0.017 (0.003) CSC Unadjusted 0.459 (0.037) 0.095 (0.055) 0.054 (0.027) 0.046 (0.011) 0.076 (0.020) 0.019 (0.005) CMHE ($K=5$) 0.306 (0.016) 0.091 (0.036) 0.104 (0.012) 0.015 (0.002) 0.093 (0.011) 0.232 (0.013) CMHE ($L=5$) 0.256 (0.018) 0.135 (0.020) 0.112 (0.008) 0.020 (0.003) 0.221 (0.009) 0.349 (0.006) CMHE ($K=L$) 0.372 (0.018) 0.101 (0.023) 0.181 (0.013) 0.072 (0.003) 0.119 (0.010) 0.323 (0.010) KMeans + TE 0.156 (0.012) 0.075 (0.009) 0.184 (0.016) 0.072 (0.012) 0.147 (0.014) 0.039 (0.010) Virtual Twins 0.284 (0.081) 0.034 (0.029) 0.066 (0.026) 0.029 (0.012) 0.082 (0.023) 0.036 (0.010) Observational CSC 0.446 (0.016) 0.119 (0.072) 0.087 (0.018) 0.026 (0.008) 0.097 (0.023) 0.027 (0.016) CSC Unadjusted 0.444 (0.048) 0.077 (0.046) 0.065 (0.026) 0.045 (0.020) 0.118 (0.026) 0.034 (0.022) CMHE ($K=5$) 0.303 (0.010) 0.095 (0.030) 0.112 (0.017) 0.018 (0.005) 0.094 (0.013) 0.229 (0.010) CMHE ($L=5$) 0.217 (0.026) 1.341 (2.717) 0.119 (0.014) 0.022 (0.005) 1.432 (2.720) 1.561 (2.709) CMHE ($K=L$) 0.371 (0.029) 0.105 (0.009) 0.208 (0.043) 0.076 (0.006) 0.186 (0.025) 0.325 (0.004) KMeans + TE 0.162 (0.011) 0.058 (0.005) 0.210 (0.031) 0.084 (0.018) 0.122 (0.032) 0.032 (0.006) Virtual Twins 0.320 (0.084) 0.062 (0.038) 0.026 (0.008) 0.044 (0.021) 0.077 (0.040) 0.038 (0.018)
+Table 4: Cross-validated performance (with standard deviation in parenthesis) when $K=5$.
+B.4.4 Impact of treatment rate
+Section B.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.
+Table 5 presents the performance under these different treatment rates in an observational setting.
+These results highlight CSC’s capacity to identify the underlying subgroups with the highest Rand-Index and one of the best cluster treatment effect recovery in these settings.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.790 (0.036) 0.036 (0.009) 0.034 (0.009) 0.018 (0.006) CSC Unadjusted 0.849 (0.033) 0.040 (0.005) 0.019 (0.005) 0.014 (0.004) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.149 (0.080) 0.188 (0.004) - 0.154 (0.005) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.808 (0.043) 0.035 (0.012) 0.027 (0.006) 0.019 (0.007) CSC Unadjusted 0.687 (0.048) 0.045 (0.011) 0.035 (0.016) 0.028 (0.006) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.010) 0.077 (0.005) 0.069 (0.004) CMHE ($L=3$) 0.318 (0.140) - - 0.132 (0.008) CMHE ($K=L$) 0.483 (0.104) 0.208 (0.013) - - KMeans + TE 0.002 (0.003) 0.195 (0.015) 0.079 (0.012) 0.148 (0.011) Virtual Twins 0.570 (0.013) 0.030 (0.005) 0.024 (0.013) 0.022 (0.005)
+Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates.
+B.4.5 Aligned covariates and treatment effect structure
+Section B.1 presents a data generation with the last 8 covariates having a clustered structures independent of the treatment response.
+This section explores a setting where all clusters in the covariates are associated with different treatment responses of interest.
+Specifically, we sample the last dimensions from standard normal distributions instead of various clusters, i.e.,
+$X_{[3-10]}\sim\text{MVN}(0,I^{8})$
+Table 6 summarises the performance in this setting, evidencing an improvement of the Kmeans+TE approach.
+When covariates’ clusters are aligned with the outcome of interest, this method recovers the clustering structure well, as shown by the best Rand-Index.
+Non-randomisation of treatment, however, worsens the treatment estimates due to the absence of correction in the estimation of treatment responses.
+Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses.
+This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across settings.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.487 (0.052) 0.027 (0.013) 0.047 (0.020) 0.030 (0.009) CSC Unadjusted 0.522 (0.041) 0.025 (0.004) 0.027 (0.018) 0.034 (0.013) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.289 (0.078) 0.063 (0.009) 0.050 (0.019) 0.146 (0.020) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.666 (0.146) 0.016 (0.007) 0.029 (0.013) 0.033 (0.010) CSC Unadjusted 0.613 (0.207) 0.024 (0.016) 0.039 (0.015) 0.030 (0.003) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.294 (0.123) 0.057 (0.005) 0.061 (0.007) 0.144 (0.007) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
+Table 6: Cross-validated performance (with standard deviation in parenthesis) when all covariates clusters presents different treatment responses.
+This setting aligns with the assumptions made by the step-wise approaches.
+B.4.6 Linear treatment responses
+As an alternative setting, we propose a setting with a linear treatment response, as assumed by CMHE Nagpal et al.
+(2023).
+To this end, we followed a similar process as described in Section B.1 at the exception of the treatment response.
+We draw the response under no treatment from a Gompertz, and the one under treatment with the same parameters shifted by the treatment effect.
+Formally, using the same notation
+$B^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
+$\Gamma^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
+$T_{0}\mid Z,X,B^{0}_{z},\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
+$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{0}_{z},A_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0},a_{z})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x)\times a_{z},s_{0}(\gamma_{z}^{0},x)\right)$
+with $a_{z}$ the treatment effect for cluster $z$. In this experiments, we choose $a_{0},a_{1},a_{2}=(0.5,1,1.5)$.
+Table 7 summarises performances in this scenario.
+While demonstrating a significant improvement of the CMHE alternatives, CSC presents competing performance on subgroup identification and treatment effect estimates.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.389 (0.065) 0.033 (0.015) 0.058 (0.035) 0.014 (0.006) CSC Unadjusted 0.411 (0.125) 0.029 (0.017) 0.062 (0.021) 0.017 (0.015) CMHE ($K=3$) 0.304 (0.033) 0.040 (0.003) 0.050 (0.003) 0.030 (0.003) CMHE ($L=3$) 0.211 (0.109) 0.051 (0.003) 0.106 (0.006) 0.028 (0.008) CMHE ($K=L$) 0.384 (0.060) 0.411 (0.797) - 0.435 (0.650) KMeans + TE 0.101 (0.139) 0.080 (0.011) 0.064 (0.028) 0.038 (0.023) Virtual Twins 0.265 (0.085) 0.025 (0.006) 0.018 (0.004) 0.064 (0.013) Observational CSC 0.311 (0.043) 0.029 (0.003) 0.074 (0.028) 0.027 (0.016) CSC Unadjusted 0.275 (0.168) 0.032 (0.005) 0.044 (0.044) 0.026 (0.014) CMHE ($K=3$) 0.360 (0.068) 0.038 (0.002) 0.050 (0.002) 0.035 (0.009) CMHE ($L=3$) 0.285 (0.111) 0.055 (0.006) 0.094 (0.007) 0.017 (0.009) CMHE ($K=L$) 0.382 (0.058) 0.052 (0.011) 0.128 (0.013) 0.028 (0.020) KMeans + TE 0.115 (0.154) 0.068 (0.020) 0.076 (0.031) 0.031 (0.008) Virtual Twins 0.186 (0.117) 0.048 (0.027) 0.041 (0.010) 0.039 (0.012)
+Table 7: Cross-validated performance (with standard deviation in parenthesis) when considering a linear treatment response. This setting aligns with the assumption made by CMHE.
+Appendix C Real-world case Study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
 To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017.
-Following [25, 10, 20], we select women who died from the condition or from cardiovascular diseases.
+Following Lee et al.
+(2018); Danks & Yau (2022); Jeanselme et al.
+(2023), we select women who died from the condition or from cardiovascular diseases.
 From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes.
 To this end, we subselect patients with recorded treatment and who received chemotherapy.
-These criteria led to the selection of 239,855 patients with 22 tumour characteristics measured at diagnosis.
+These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  Danks & Yau (2022) for further description).
 Selection of $K$.
@@
 In the absence of experts’ intuition, one may rely on an elbow rule heuristic over $l_{F}$.
-This means selecting the number of subgroups after which one observes limited likelihood improvement as used in [19].
 Using Figure 6, the negative log-likelihood presents an elbow for $K=2$.
@@
 In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy.
-This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist [24], making this setting more likely to meet the positivity assumption (Assumption 6), necessary to study causality in observational data.
+This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist Lazzari et al.
+(2023), making this setting more likely to meet the positivity assumption (assumption 6), necessary to study causality in observational data.
 (a) Neural Survival Treatment
-(b) Cox Mixture of Heterogenous Effects
+(b) Cox Mixture of Heterogeneous Effects
 Figure 7: Averaged treatment effect subgroups across 5-fold cross-validation observed in the Seer dataset with the shaded areas representing 95% CI.
 Uncovering treatment response.
-Figure 6 presents the identified treatment effect subgroups when using CNSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning.
+Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning.
 As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups.
@@
 Figure 8 displays the 10 covariates most indicative of the different treatment response subgroups.
-Figure 7: Causal Neural Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
-Table 2 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) [42].
+Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
+Table 8 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) Royston & Parmar (2013).
 Both methodologies identify a population with limited treatment response.
@@
 RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.00 (0.00) 93.3% 55.5% 1.18 (5.74) 17.4% 46.6% Subgroup 1 0.62 (0.16) 6.7% 47.5% 17.75 (16.08) 23.4% 48.6%
-Table 2: Causal Neural Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
+Table 8: Causal Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
 The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy.
 However, our methodologies remain hypothesis-generating tools, requiring further experimental validation, particularly due to potential confounding through hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options.
+The quality of available covariates limits this analysis and only serves as an example to medical practitioners to identify subgroups of treatment responses from observational data.

### v4

@@
 Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials.
-Existing approaches for subgroup analysis primarily rely on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised.
-RCTs’ patient cohorts are often constrained by cost, rendering them not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice.
-When applied to observational studies, subgroup analysis approaches suffer from significant statistical biases particularly because of the non-randomisation of treatment.
-Our work introduces a novel, outcome-guided method for identifying treatment response subgroups in observational studies.
-Our approach assigns each patient to a subgroup associated with two time-to-event distributions: one under treatment and one under control regime.
-It hence positions itself in between individualised and average treatment effect estimation.
-The assumptions of our model result in a simple correction of the statistical bias from treatment non-randomisation through inverse propensity weighting.
-In experiments, our approach significantly outperforms the current state-of-the-art method for outcome-guided subgroup analysis in both randomised and observational treatment regimes.
+Existing approaches for treatment effect estimation primarily rely on Randomised Controlled Trials (RCTs), which are often limited by insufficient power, multiple comparisons, and unbalanced covariates.
+In addition, RCTs tend to feature more homogeneous patient groups, making them less relevant for uncovering subgroups in the population encountered in real-world clinical practice.
+Subgroup analyses established for RCTs suffer from significant statistical biases when applied to observational studies, which benefit from larger and more representative populations.
+Our work introduces a novel, outcome-guided, subgroup analysis strategy for identifying subgroups of treatment response in both RCTs and observational studies alike.
+It hence positions itself in-between individualised and average treatment effect estimation to uncover patient subgroups with distinct treatment responses, critical for actionable insights that may influence treatment guidelines.
+In experiments, our approach significantly outperforms the current state-of-the-art method for subgroup analysis in both randomised and observational treatment regimes.
+Machine Learning, ICML
 1 Introduction
-Understanding heterogeneous therapeutic responses between patient subgroups is the core of treatment guidelines and the development of new drugs.
-Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed (Foster et al., 2011).
-For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes (Investigators, 1996), shaping subsequent guidelines.
-Figure 1 illustrates this idea: two groups may present opposite treatment responses, and would benefit from different recommendations.
-Our work aims to uncover such subgroups.
-Randomised controlled trials (RCTs) remain the gold standard for identifying subgroups of treatment effects.
-In RCTs, patients are randomly assigned to control or treated groups, allowing researchers to assess the impact of an intervention or treatment.
-However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies (Hernán & Robins, 2016).
-Our work introduces a novel approach that diverges from traditional RCT-based methodologies (Nagpal et al., 2022a; 2023) by leveraging routinely collected observational data to identify patient subgroups.
-Observational studies encompass larger and more diverse cohorts reflective of real-world practices, offering the potential to uncover subgroups of treatment responses, that could be missed in RCTs.
-Prior works (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017) have leveraged observational data while addressing biases from non-random treatment assignments (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018).
-However, this body of literature primarily focuses on estimating (i) averaged treatment effects at the population level or (ii) individualised treatment effects, thereby overlooking the identification of treatment effect subgroups.
-Addressing this critical gap in the literature, our work uncovers patient subgroups with distinct treatment responses using observational data.
-Unlike previous methods relying on RCTs, we introduce a mixture of neural networks as an extension to traditional outcome-guided models to uncover subgroups delineated by non-linear, higher-dimensional combinations of available covariates, without requiring parametrisation of the time-to-event distributions or treatment responses.
+Understanding heterogeneous therapeutic responses among patient subgroups is central to developing clinical guidelines and new treatments.
+Identifying such subgroups is valuable to inform the design of future clinical trials and to direct healthcare resources to those most likely to benefit, and away from those at greatest risk of harm (Foster et al., 2011).
+A striking example comes from the BARI trial on coronary artery disease, which suggested that patients with diabetes should receive coronary artery bypass grafts rather than percutaneous interventions, whereas patients without diabetes benefited from the opposite strategy (Investigators, 1996), an observation that shaped subsequent guidelines.
+Figure 1 illustrates this concept: two groups with opposing treatment responses may require distinct therapeutic recommendations.
+Our work aims to uncover such subgroups in observational, time-to-event data.
+Randomised controlled trials (RCTs) remain the gold standard for studying heterogeneous treatment effects.
+By randomly assigning patients to control or treated arms, RCTs eliminate potential confounders when assessing a treatment’s impact on outcomes.
+However, RCTs are often costly, time-consuming, and restricted to specific patient cohorts, which are likely not to reflect the full spectrum of patients seen in clinical practice (Hernán & Robins, 2016, 2010; Cole & Hernán, 2008).
+Trial findings may thus not generalise to real-world settings, where treatment choice depends on clinical judgment and patient preferences, among other variables that may remain unobserved.
 Figure 1: Subgroup treatment effect discovery in time-to-event observational data.
-Our method identifies subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials.
-Our method simultaneously models the treatment effect and identifies subgroup while addressing censoring and treatment non-randomisation.
-Contributions.
-Our work introduces the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects.
-Section 2 formalises the problem of treatment subgroup discovery and non-randomisation correction.
-Then, we introduce the proposed monotonic neural network implementation in Section 2.
-Finally, we evaluate the proposed methodology on two synthetic datasets in Section 3, an extensive set of alternative settings in App.
+Our work aims to identify subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials.
+Our method simultaneously models the treatment effect and identifies subgroup while addressing treatment non-randomisation and censoring.
+To address these limitations, researchers increasingly leverage large-scale observational datasets which capture diverse, real-world populations.
+Although observational data offer the potential to identify subgroups that RCTs might miss, they also pose challenges due to non-random treatment assignments and confounding (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018).
+A robust methodological literature in causal inference has emerged to tackle these challenges.
+Techniques such as inverse probability weighting (Cole & Hernán, 2008), marginal structural models (Robins et al., 2000), and doubly robust estimators (Bang & Robins, 2005) allow for more accurate causal effect estimation under appropriate assumptions.
+Prior works in machine learning (ML) have extended these ideas to estimate treatment effects from observational data (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017).
+However, most methods focus on (i) average treatment effects at the population level or (ii) individualised treatment effects.
+These approaches overlook the utility of identifying treatment-effect subgroups —groups with consistently distinct responses to a given intervention.
+Detecting such subgroups is crucial in translating treatment effect estimates into actionable insights for clinical decision-making and resource allocation.
+We address this gap by introducing a novel framework that uncovers patient subgroups with distinct treatment responses using observational time-to-event data.
+Extending the rich tradition of outcome-guided modelling (Foster et al., 2011) to a flexible, neural network–based architecture.
+Crucially, our approach does not require explicit parametrisation of time-to-event distributions or treatment effects; it instead leverages a mixture of monotonic neural networks with correction for non-random treatment assignment.
+In summary, our contributions are:
+  • Neural network-based treatment subgroup identification.
+Our work introduces the first neural network architecture to simultaneously estimate and identify subgroups of treatment effects in observational settings, accommodating both censoring and non-random treatment assignments without parameterising the treatment effects.
+  • Novel formalisation.
+In section 2, we review existing approaches to identify subgroups of treatment effects.
+Section 3 formalises the problem of subgroup discovery and our strategy for minimising biases arising from non-random treatment assignment.
+  • Extensive application on survival time-to-event data.
+In section 3, we introduce the proposed monotonic neural network implementation, and evaluate its performance on two synthetic datasets in Section 4, an extensive set of alternative settings in App.
 B and a real-world example described in App.
-C and compare it with related approaches reviewed in Section 4.
-2 Method
-2.1 Problem setup
-Latent survival subgroup analysis.
-Our goal is to uncover subgroups with similar treatment responses, guided by the observed times of occurrence of an outcome of interest.
-Patients are assigned to subgroups based on their covariates at treatment time.
-Each subgroup is characterised by two distributions, known as survival functions: one under treatment and one under control regimes.
-These distributions characterise the probability of observing the event of interest after a given time under a given treatment.
-Consider the random variables associated with observed covariates $X$, an indicator identifying whether the event of interest, in our analysis death, was observed $D$, and the observed time of event $T$.
-Formally, we define the latter random variables as $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is random variable of the (right)-censoring time, i.e.
-the time at which the patient left the study before experiencing the event of interest, and $T^{\prime}$ is the partially observed random variable associated with the time of the event of interest if there was no censoring.
-When a patient is censored, $T=C$ and $D=0$, otherwise the event is observed and $T=T^{\prime}$ and $D=1$.
-To estimate the associated survival likelihood, we further assume, as commonly done in survival analysis, that the censoring time $C$ is not informative for the time of the event of interest $T^{\prime}$.
-Assumption 1 (Non-informative censoring).
-The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$ given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
-Central to our problem is the additional variable $Z$ associated with the latent, unobserved subgroup membership.
-Following Jeanselme et al.
-(2022), we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3..
-We assume that $Z$ depends on $X$ and influences $T^{\prime}$.
-As we are interested in recovering the survival functions associated with the latent subgroups, we ignore the potential dependence of $T^{\prime}$ on $X$, considering the individual survival distribution as a mixture of the different subgroups.
-Assumption 2 (Mixture modelling).
-The event time $T^{\prime}$ is completely determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
-While this assumption may hurt individual performance as individual covariates do not directly inform individual survival but group membership only, it improves the interpretability of the model by separating subgroup membership from survival profiles.
-In conclusion, we aim to recover $Z$ and the survival distributions associated with each subgroup from the observed $X$, $T$ and $D$.
-Given the previously described dependencies, the expected survival $S$ for a given patient with covariates $x$ is the survival when marginalising over the different subgroups:
-$S(t\mid X=x):=\mathbb{P}(T^{\prime}\geq t\mid X=x)=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid X=x)\mathbb{P}(T^{\prime}\geq t\mid Z=k)$ (1)
-Note that the last term is linked to the quantity commonly used in the survival literature: $\Lambda_{k}(t):=-\log\mathbb{P}(T^{\prime}\geq t\mid Z=k)$, the group-specific cumulative hazard, with derivative with respect to time $\lambda_{k}(t)$, the group-specific instantaneous hazard, corresponding to the increase in the probability of observing the event of interest given survival until that time $t$.
-Estimating the two probability distributions $\mathbb{P}(Z=k\mid X=x)$ and $\mathbb{P}(T^{\prime}\geq t\mid Z=k)$ is our core interest.
-Likelihood-based model fitting.
-One can maximise the log-likelihood associated with the observed realisations $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i\in[\![1,N]\!]$ with $N$ the total number of patients (we omit $i$ where redundant) to estimate the survival function (Eq.
-equation 1).
-Under Assumption 1, Equation equation 2 describes this factual [^2] Referring to the likelihood of the observed (rather than counterfactual) realisations.
-log-likelihood.
-$l_{F}(\theta)$ $=\sum_{i,d_{i}=1}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}\log S_{\theta}(t_{i}\mid x_{i})$ (2)
-where $\theta$ is the set of parameters characterising the estimated survival function $S_{\theta}$.
-Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see Jeanselme et al.
-(2022) for derivation).
-Each censored patient ($d_{i}=0$) contributes the probability of not observing an event before $t_{i}$, i.e., $S_{\theta}(t_{i}\mid x_{i})$, under assumption 1 of independence between the censoring time and the event of interest conditional on the covariates.
-Latent treatment effect subgroups.
-Consider now the binary treatment variable $A$.
-Patients either receive the treatment ($A=1$) or they do not ($A=0$).
-Therefore, following the potential outcomes formulation, we consider the time of events under treatment $T^{\prime}_{1}$ and under the control regime $T^{\prime}_{0}$.
-The central challenge is that one can not observe both $T^{\prime}_{0}$ and $T^{\prime}_{1}$, but only $T^{\prime}$, the random variable associated with the event time under the observed treatment regime in the absence of censoring, which is equal to $A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under assumption 4 described below.
-Figure 2 describes the dependencies between the previously described random variables.
-A critical assumption of our proposed setting is that the subgroups we aim to identify do not influence treatment if we know a patient’s covariates, formalised as follows:
-Assumption 3 (Unknown latent groups).
-The treatment assignment $A$ is independent of the subgroup membership $Z$ given the covariates $X$. Formally, $A\perp\!\!\!\perp Z\mid X$.
-This assumption relaxes the traditional assumption of treatment randomisation.
-Note that this assumption remains realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) if subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
-Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
-Further, the following three assumptions are necessary to estimate treatment effects and are common in the causal literature.
-Assumption 4 (Consistency).
-A patient’s observed event time is the potential event time associated with the observed treatment.
-Formally, this means $T^{\prime}:=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the treatment $A$.
-Assumption 5 (Ignorability).
-The potential event times are independent of the treatment given the observed covariates, i.e.
-$A\perp\!\!\!\perp(T^{\prime}_{0},T^{\prime}_{1})\mid X$.
-Equivalently, no unobserved confounders impact both treatment and event time.
-Assumption 6 (Overlap / Positivity).
-Each patient has a non-zero probability of receiving the treatment, i.e. $\mathbb{P}(A\mid X)\in(0,1)$ where $(0,1)$ is the open interval, resulting in a non-deterministic treatment assignment.
-Under the Assumptions 4, 5, and 6, existing works often focus on estimating the Individualised Treatment Effect (ITE, here denoted as $\tau$), defined as the difference between survival under the two potential treatment regimes given the covariates $x$ (see App.
-A.1 for derivation):
-$\tau(t,x)=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (3)
-Estimating this quantity requires accurately modelling the survival distributions under the two treatment regimes for all $x$ from observed data.
-This estimation would be straightforward if one could access the observed survival times for all patients under both treatment regimes.
-In this case, one would estimate the survival using the event times under $A=1$ and under $A=0$.
-The key challenge is that the counterfactual survival outcome is not observed: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
-In RCTs, treatment $A$ and covariates $X$ are by design independent.
-Hence, the subset of the patients receiving either treatment regime is representative of the overall population.
-Relying on this observed subset to estimate the survival distribution under each regime is a valid estimate for the population.
-In other words, maximising the factual likelihood to estimate survival under both treatment regimes results in a valid estimate.
-However, covariates and treatment are not independent in observational studies in which treatment recommendations depend upon the observed covariates.
-Formally, $\mathbb{P}(A\mid X)\neq\mathbb{P}(A)$ in general.
-For example, clinicians might recommend more aggressive treatment to patients in more severe conditions.
-This absence of randomisation results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021).
-In this setting, estimating survival by maximising the factual likelihood is no-longer enough to estimate the counterfactual survival distribution.
-Under assumption 6, approaches such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity of learnt representations (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Shalit et al., 2017) are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest.
-Specifically, Shalit et al.
-(2017) demonstrate that the negative log likelihood is upper-bounded by observable quantities.
-Here we extend their notations to the survival setting as
-$-l(\theta)\leq-l_{F}^{*}(\theta)+\gamma\cdot\text{IPM}(q_{\Phi}^{A=0},q_{\Phi}^{A=1}),$ (4)
-where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in equation 2) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
-Shalit et al.
-propose to use this bound to train a neural network to estimate treatment effects.
-In this context, $\Phi$ is an inner representation of the network and the IPM regularisation renders this embedding similar between the treated and untreated populations.
-This embedding is then used to estimate the survival function under both treatment regimes.
-The difference between the estimated survival functions is then an accurate estimate of the treatment effect as the IPM penalisation corrects for the shift resulting from treatment non-randomisation.
-Note that, under the dependencies we assume, only the subgroup assignment ($\mathbb{P}(Z\mid X)$) depends on the covariates $X$ and could, consequently, play the role of the transformation $\Phi$.
-However, treatment is assumed independent of the group membership under assumption 3, meaning that treatment rate does not differ between subgroups.
-This assumption results in the regularisation term being null, i.e.
-$\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$ (see App.
-A.2 for derivation).
-This assumption is consequently necessary to correctly estimate the subgroups’ treatment effect, without penalising the clustering structure.
-As a consequence, our work focuses on the first term of the upper bound in equation 4 by re-weighting the factual likelihood with the patient’s propensity score (Hassanpour & Greiner, 2019; Shalit et al., 2017).
-Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting (Austin & Stuart, 2015) scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
-$w_{i}^{-1}=\begin{cases}0.05&\text{ if }\hat{p}_{A}(x_{i})<0.05\\ 0.95&\text{ if }\hat{p}_{A}(x_{i})>0.95\\ a_{i}\cdot\hat{p}_{A}(x_{i})+(1-a_{i})\cdot(1-\hat{p}_{A}(x_{i}))&\text{ otherwise},\end{cases}$ (5)
-where $i$ is the patient index, and $a_{i}$ is the realisation of $A$ for patient $i$.
-Using the factual likelihood from equation 2 and the weights $w_{i}$ defined in equation 5, we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
-$\sum_{i,d_{i}=1}w_{i}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i},a_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}w_{i}\log S_{\theta}(t_{i}\mid x_{i},a_{i})$ (6)
-Figure 3: Causal Survival Clustering architecture.
-Latent parameter $l_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the cumulative hazard $\Lambda_{k}$ under both treatment regimes.
-$G$ assigns the probability to belong to each subgroup given the patient’s covariate(s) $x$.
-To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood.
-In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
-2.2 Estimating the quantities of interest with neural networks
-The previous section discussed the quantities one must estimate —here parameterised by neural networks— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the cumulative incidence $\Lambda_{k}$ characterising survival under the two treatment regimes, and the weights $w$.
-Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
-Subgroup assignment.
-Similar to Jeanselme et al.
-(2022), a multi-layer perceptron $G$ with final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
-$G(x):=[\mathbb{P}(Z=k\mid x)]_{k=1}^{K}$
-Survival distributions.
-Each subgroup $k$ is represented by a vector $l_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the cumulative hazard functions.
-The vector $l_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic positive outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in Jeanselme et al.
-(2022; 2023)., with a final SoftPlus layer to model the cumulative hazard functions under both treatment regimes $\Lambda_{k}(t):=(\Lambda_{k}(t\mid A=0),\Lambda_{k}(t\mid A=1))$.
-We use the following transformation to ensure that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
-$\Lambda_{k}(t):=t\cdot M(l_{k},t)$
-By modelling these two cumulative intensity functions, we can estimate the Subgroup Average Treatment Effect, $\hat{\tau}_{k}(t)$, in subgroup $k$ as
-$\hat{\tau}_{k}(t):=\mathbb{E}[\tau(t,X)\mid Z=k)]=\exp(-\Lambda_{k}(t\mid A=1))-\exp(-\Lambda_{k}(t\mid A=0))$
-Inverse propensity weighting.
-Under treatment randomisation, as in RCTs, the previous components would accurately model the observed survival outcomes and identify subgroups of treatment effects by maximising the factual likelihood.
-As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity score of a patient estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as
-$W(x):=\mathbb{P}(A=1\mid x)$
-2.3 Training procedure
-First, the network $W$ is trained to predict the binary treatment assignment by minimising the cross-entropy of receiving treatment.
-Then, training all other components relies on minimising the weighted factual log-likelihood introduced in equation 6.
-The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as automatic differentiation of the monotonic neural networks’ outcomes readily provides the instantaneous hazards $\lambda_{k}$ necessary for computing the survival function derivative in the log likelihood (Jeanselme et al., 2022; 2023; Rindt et al., 2022).
-3 Experimental analysis
-As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known.
-Appendix C accompanies these synthetic results with the analysis of heterogeneity in real-world, adjuvant radiotherapy responses for patients diagnosed with breast cancer.
-Our code to produce the synthetic dataset, the model, and the reproduction of all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
-3.1 Data generation
-We generate a population of $3,000$ equally divided into $K=3$ subgroups.
-We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parameterised by group membership and individual covariates.
-Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model.
-This choice reflects a setting in which time-to-event distributions under treatment or control regimes are different functions of the covariates.
-This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis which often assume a linear treatment response, i.e., a shift between treated and untreated distributions.
-Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational).
-Finally, non-informative censoring times are drawn following a Gompertz distribution.
-Details of the generative process of our synthetic dataset are deferred to App.
-B.1.
-3.2 Empirical settings
-Benchmark methods.
-We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE Nagpal et al.
-(2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups.
-This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted.
-The central differences with our proposed approach is that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response.
-This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival.
-Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses.
-By contrast, our approach identifies subgroups of treatment while considering survival, without constraining the treatment response.
-We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, in comparison to a group with the same treatment responses but better survival odds.
-Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
-For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups.
-Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
-Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
-Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately.
-First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in Nagpal et al.
-(2022a), referred as Kmeans + TE in the following.
-Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters.
-However, this approach is limited by its implicit assumption of randomisation and its pre-specified number of subgroups, which is not informed by observed outcomes.
-Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them.
-As proposed in the literature, we propose to use a survival tree to estimate response under each treatment regime and then use a KMeans on the estimated treatment effects, computed as the difference between survival estimates.
-Note that this methodology also assumes treatment randomisation, and the clustering is uninformed by the covariates.
-For details on training and hyperparameter optimisation, we refer to App.
-B.2.
-Evaluation.
-In the synthetic experiments, the subgroup structure is known.
-We measure the adjusted [^6] Random patient assignment results in an adjusted Rand-Index of 0.
-Rand-Index (Rand, 1971), which quantifies how the estimated assignment aligns with the known underlying group structure.
-Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
-[…]
-where $\hat{\tau}_{\hat{k}}$ is the estimated treatment effect for subgroup $\hat{k}=\operatorname*{arg\,max}_{l}\mathbb{E}_{x\in k}(Z=l\mid x)$, i.e.
-the most likely assigned cluster for patients in the underlying $k$ cluster, and $\tau_{k}$ is the ground truth.
-3.3 Treatment effect recovery
-In this section, we present the results under the previously described data generation process. App. B.4 presents alternative datasets which demonstrate the flexibility of the proposed methodology.
-Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the "Observational" treatment assignment with the shaded area representing 95% CI.
-The log-likelihood presents an elbow around the underlying number of subgroups.
-Recovering the underlying number of subgroups.
-For all methodologies, we need to choose the number of subgroups a priori.
-An important question is, therefore, whether we can identify the true number of subgroups with our model.
-Figure 4 presents the average NLL obtained by cross-validation for models with different numbers of subgroups $K$.
-The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood.
-Using this heuristic, the optimal choice for $K$ is $3$, which agrees with the underlying generative process.
-App.
-B.3 explores the sensitivity to this choice, demonstrating the method’s robustness.
-This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses which separate clustering from treatment effect estimates.
-In these methods, survival outcomes cannot directly guide the choice of $K$.
-Discovering subgroups.
-Next, we present quantitative results of our method against the benchmarks methods.
-Table 1 presents the performance of the different methodologies under the two studied scenarios.
-Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival distributions.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.798 (0.054) 0.042 (0.012) 0.032 (0.007) 0.015 (0.006) CSC Unadjusted 0.804 (0.052) 0.032 (0.009) 0.027 (0.007) 0.018 (0.010) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.258 (0.155) - 0.088 (0.007) - CMHE ($K=L$) 0.392 (0.054) 0.338 (0.262) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.812 (0.038) 0.037 (0.013) 0.034 (0.009) 0.022 (0.005) CSC Unadjusted 0.727 (0.084) 0.048 (0.019) 0.025 (0.011) 0.025 (0.009) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
-Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation.
-Best performance per column and simulation scenario are marked in bold, second best in italic.
-’-’ describes when the methodology diverges.
-Our proposed CSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
-CSC outperforms all CMHE alternative, the current state-of-the-art method.
-CMHE’s parametrisation and assumptions explain this difference.
-Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log-hazards.
-Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data.
-In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes.
-Increasing the number of subgroups, as shown in $K=L$ improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods.
-These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
-CSC presents the best performance in identifying the underlying subgroup.
-Our proposed CSC method outperforms all approaches on the Rand-Index.
-In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are aligned.
-KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index.
-Consequently, the identified clusters differ in their treatment responses as shown by the large IAE.
-The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE.
-However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates.
-This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
-Treatment assignment correction improves subgroup identification in observational settings.
-The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation.
-The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context.
-Note that estimating propensity slightly reduces performance due to error introduced, reducing performance for the corrected alternative.
-In a randomised setting, CSC Unadjusted should be preferred.
-However, the superiority of CSC over its unadjusted alternative is evident in the observational setting, where CSC better recovers the different groups as shown by the Rand-Index and subgroup treatment effects.
-Finally, Appendices B.4.1 to B.4.5 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
-Real-world heterogeneity analysis in adjuvant radiotherapy responses (App. C).
-Using the Surveillance, Epidemiology, and End Results program [^7] Available at https://seer.cancer.gov/(Seer), App.
-C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research.
-Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment.
-Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis.
-Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
+C, benchmarking against several established approaches.
 2 Related work
-While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a; b; Jeanselme et al., 2022) to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse.
-The current literature focuses on estimating population or individual treatment effects (Curth et al., 2021; Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020).
-While these models help to understand the average population-wide response to treatment and aim to estimate individualised treatment effects, they do not provide an understanding of the patient groups that may benefit more or be harmed more by treatment.
+While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a, b; Jeanselme et al., 2022) to identify different phenotypes of patients, the ML literature on phenotyping treatment effects with time-to-event outcomes remains sparse.
+The current focus is on estimating population or individual treatment effects in the static (Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020) and survival settings (Curth et al., 2021).
+While these approaches help to understand the population or individualised treatment effects, they do not identify subgroups that may benefit or be harmed by treatment.
 Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
@@
 (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits.
-Athey & Imbens (2015; 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation.
+Athey & Imbens (2015, 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation.
 Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation.
-Each obtained split in the decision tree delineates two subgroups of treatment effect Lipkovich et al.
-(2011); Loh et al.
-(2015).
+Each obtained split in the decision tree delineates two subgroups of treatment effect (Lipkovich et al., 2011; Loh et al., 2015).
 Alternatively, McFowland III et al.
@@
 (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
-Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with inherently continuous survival outcomes that our paper explores.
+Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with survival outcomes that our work explores.
 At the intersection with survival analysis, Zhang et al.
@@
 As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
+3 Method
+3.1 Problem setup
+Our goal is to uncover a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 4.
+with distinct treatment responses, guided by the observed times of occurrence of an outcome of interest.
+To this end, our model assigns patients to latent subgroups based on their covariates at the treatment decision time.
+For each subgroup, we estimate the probability of not observing the event of interest given time under treatment and under control regimes.
+These two distributions are known as survival functions, and their difference corresponds to the estimated treatment effect for a given group.
+Formally, consider the random variables associated with observed covariates $X$, the assigned binary treatment $A$, and the observed event time $T^{\prime}$.
+Following the potential outcomes formulation, $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under consistency (Assumption 1) where $T^{\prime}_{1}$ is the potential event time under treatment and $T^{\prime}_{0}$, under the control regime.
+Assumption 1 (Consistency).
+A patient’s observed event time is the potential event time associated with the observed treatment.
+Formally, this means $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the control and treatment regimes respectively.
+Central to our problem is the latent, unobserved, subgroup membership $Z$.
+As formalised in Assumption 2, we assume that the variable $Z$ mitigates the dependence between the covariates $X$ and the observed event time $T^{\prime}$.
+Assumption 2 (Mixture mitigation).
+The event time $T^{\prime}$ is determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
+In this setting, our interest lies in estimating the Subgroup Average Treatment Effect (SATE) for group $k$:
+$\tau_{k}(t)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,Z=k)$
+$-\mathbb{P}(T^{\prime}\geq t\mid A=0,Z=k)$ (1)
+This focus differs from the literature interest in estimating Individualised Treatment Effects (ITE), defined as follows (see App. A.1 for derivation from the potential outcomes formalisation):
+$\tau(t,x)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)$
+$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$
+$=\sum_{k}\tau_{k}(t)\mathbb{P}(Z=k\mid X=x)$ (Under Asm. 2)
+Estimating the previous quantities requires the accurate modelling of the survival distributions under the two treatment regimes.
+The central challenge in this estimation is that the counterfactual survival outcome is unobserved: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
+The absence of treatment randomisation in observational studies, e.g.
+clinicians might recommend more aggressive treatment to patients in more severe conditions, results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021).
+This prohibits the estimation of the survival functions through the maximisation of the observed outcomes likelihood.
+Under Assumptions 3 and 4, remedies such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity between treatment regimes (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Hernán & Robins, 2006; Shalit et al., 2017) allows to estimate both factual and counterfactual likelihood from the observed data by penalising the difference between the treated and non-treated populations.
+In our work, we adopt an Inverse Propensity Weighting (IPW) of the factual likelihood to estimate the overall likelihood, consisting in weighting each observation by the inverse probability of receiving treatment.
+Assumption 3 (Ignorability).
+The potential event times are independent of the treatment given the observed covariates, i.e.
+$A\perp\!\!\!\perp(T^{\prime}_{0},T^{\prime}_{1})\mid X$.
+Equivalently, no unobserved confounders impact both treatment and event time.
+Assumption 4 (Overlap / Positivity).
+Each patient has a non-zero probability of receiving the treatment, i.e. $\mathbb{P}(A\mid X)\in(0,1)$ where $(0,1)$ is the open interval, resulting in a non-deterministic treatment assignment.
+To express the factual likelihood, one needs to tackle a final challenge in time-to-event data: censoring.
+Patients may not observe the outcome of interest during the study period or exit the study for reasons independent of the outcome of interest.
+Formally, instead of observing $T^{\prime}$, one observes an indicator identifying whether the event of interest was observed $D$, and the associated observed time of event $T$, with $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is the random variable of the (right)-censoring time.
+When a patient is censored, $T=C$ and $D=0$; otherwise, the event under treatment regime $A$ is observed and $T=T^{\prime}_{A}$ and $D=1$.
+As ignoring the censoring process biases the likelihood, we assume, as commonly done in survival analysis, that this process is uninformative, formalised as:
+Assumption 5 (Non-informative censoring).
+The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$ given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
+Under the previous assumptions, we can maximise the overall likelihood by optimising for the following weighted factual log-likelihood $l_{F}$:
+$l_{F}=$ $\sum_{i,d_{i}=1}w_{i}\log\left(\frac{\partial S(t\mid x_{i},a_{i})}{\partial t}\bigg{|}_{t=t_{i}}\right)$
+$+\sum_{i,d_{i}=0}w_{i}\log S(t_{i}\mid x_{i},a_{i})$ (2)
+where $i$ is the patient index, $(x_{i},t_{i},a_{i},d_{i})$ are realisations of the associated variables $X,T,A$ and $D$, and $w_{i}$ is the inverse propensity weighting correction for patient $i$.
+Under Assumption 2, the survival corresponds to the weighted survival over the different subgroups:
+$S(t\mid x,a):=\mathbb{P}(T^{\prime}\geq t\mid x,a)$
+$=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid x,a)\mathbb{P}(T^{\prime}\geq t\mid Z=k,a)$ (3)
+In summary, Figure 2 describes all variables and dependences assumed in the studied problem with a directed acyclic graph.
+Given the observed $X$, $A$, $T$ and $D$, our aim is to estimate the latent structure $Z$ and the associated survival distributions under the two potential treatment regimes $T^{\prime}_{0}$ and $T^{\prime}_{1}$.
+Figure 2: Graphical representation between covariates ($X$), treatment ($A$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
+Figure 3: Causal Survival Clustering (CSC) architecture.
+Latent parameter $u_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the survival under both treatment regimes.
+$G$ assigns the probability of belonging to each subgroup given the patient’s covariate(s) $x$.
+To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood.
+In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
+3.2 Estimating the quantities of interest
+The previous section discussed the quantities one must estimate —here parametrised by neural networks [^2] When interpretability is a key concern, one can consider a linear model for both subgroup assignment and IPW components, as explored in Appendix B.4.— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the distributions characterising survival under the two treatment regimes $\mathbb{P}(T^{\prime}\geq t\mid A,Z=k)$, and the IPW weights $w$.
+Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
+Subgroup assignment.
+Similar to (Jeanselme et al., 2022), a multi-layer perceptron $G$ with a final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
+$G(x):=[\mathbb{P}(Z=k\mid x)]_{k=1}^{K}$
+Survival distributions.
+Each subgroup $k$ is represented by a parameter vector $u_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the survival functions learnt through backpropagation.
+The vector $u_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in (Jeanselme et al., 2022, 2023).
+and a final SoftPlus layer to ensure positivity.
+We train this neural network to model survival through the following transformation — ensuring that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
+$\mathbb{P}(T^{\prime}_{A=a}\geq t\mid Z=k):=\exp(-t\cdot M(u_{k},t)[a])$
+Inverse propensity weighting.
+Under treatment randomisation, as in RCTs, the previous components would accurately model the survival outcomes under both treatment regimes and identify subgroups of treatment effects by maximising the factual likelihood.
+As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity of receiving treatment estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as:
+$W(x):=\mathbb{P}(A=1\mid x)$
+From this estimated probability, we further truncate the propensity score (Austin & Stuart, 2015) to avoid extreme values for the estimated weights, which would result in unstable treatment effect estimates.
+In this context, the weights $w_{i}$ are defined as:
+$\forall i,w_{i}^{\prime}:=$ $a_{i}W(x_{i})+(1-a_{i})(1-W(x_{i}))$
+$w_{i}:=$ $\max(0.05,\min(w_{i}^{\prime},0.95))$ (4)
+3.3 Training procedure
+First, the model $W$ is trained to predict the binary treatment assignment probability by minimising the cross-entropy of receiving treatment.
+Then, training all other components relies on maximising the weighted log-likelihood introduced in Equation (2).
+The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as a forward path provides the estimated survival and automatic differentiation of the monotonic neural networks’ outcomes readily provides the derivative necessary for computing the log-likelihood (Jeanselme et al., 2022, 2023; Rindt et al., 2022).
+4 Experimental analysis
+As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known.
+As a real-world case study, Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer.
+Our code to produce the synthetic dataset, the model, and all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
+4.1 Data generation
+We generate a population of $3,000$ equally divided into $K=3$ subgroups.
+We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parametrised by group membership and individual covariates.
+Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model (Assumption 2).
+This choice reflects a setting in which survival distributions under treatment or control regimes are different functions of the covariates.
+This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis, which often assume a linear treatment response.
+Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational).
+Finally, non-informative censoring times are drawn following a Gompertz distribution.
+Details of the generative process of our synthetic dataset are deferred to Appendix B.1.
+Further, Appendix B.5 explores alternative datasets to demonstrate the flexibility of the proposed methodology.
+4.2 Empirical settings
+Benchmark methods.
+We compare the proposed approach (CSC) against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE (Nagpal et al., 2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups.
+This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted.
+The central differences with our proposed approach are that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response.
+This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival.
+Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses.
+By contrast, our approach identifies treatment subgroups while considering survival without constraining the treatment response.
+We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, compared to a group with the same treatment responses but with better survival odds.
+Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
+For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups.
+Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
+Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
+Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately.
+First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in (Nagpal et al., 2022a), referred to as Kmeans + TE in the following.
+Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data, and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters.
+Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them.
+As proposed in the literature, we use a survival tree to estimate response under each treatment regime and then use a KMeans (MacQueen et al., 1967) on the estimated treatment effects, computed as the difference between survival estimates.
+Note that these methodologies assume treatment randomisation and the clustering is uninformed by the covariates.
+For details on training and hyperparameter optimisation, we refer to App.
+B.2.
+Evaluation.
+In the synthetic experiments, the subgroup structure is known.
+We measure the adjusted [^6] Random patient assignment results in an adjusted Rand-Index of 0.
+Rand-Index (Rand, 1971), which quantifies how the estimated assignment aligns with the known underlying group structure.
+Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
+[…]
+where $\hat{\tau}_{\hat{k}}$ is the estimated treatment effect for subgroup $\hat{k}=\operatorname*{arg\,max}_{l}\mathbb{E}_{x\in k}(Z=l\mid x)$, i.e.
+the most likely assigned cluster for patients in the underlying $k$ cluster, and $\tau_{k}$ is the ground truth.
+4.3 Treatment effect recovery
+Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the ”Observational” treatment assignment with the shaded area representing 95% CI.
+The log-likelihood presents an elbow around the underlying number of subgroups.
+Recovering the underlying number of subgroups.
+For all methodologies, we must choose the number of subgroups a priori.
+An important question is, therefore, whether we can identify the true number of subgroups with our model.
+Figure 4 presents the average negative log-likelihood obtained by cross-validation for models with different numbers of subgroups $K$.
+The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood.
+Using this heuristic, the optimal choice for $K$ is $3$, which aligns with the underlying generative process.
+App.
+B.3 shows that when the number of clusters is misspecified, the estimated treatment effects are unstable, the variance associated with estimated groups is hence an additional tool to select the number of clusters.
+This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which separate clustering from treatment effect estimates.
+In these methods, survival outcomes cannot directly guide the choice of $K$, whereas it is directly reflected in the proposed methodology’s performance.
+Discovering subgroups.
+Next, we present the quantitative results of our method against the benchmark methods.
+Table 1 summarises the performance of the different methodologies under the two studied scenarios.
+Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival clusters.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.807 (0.055) 0.037 (0.008) 0.023 (0.008) 0.015 (0.003) CSC Unadjusted 0.802 (0.050) 0.039 (0.011) 0.021 (0.009) 0.014 (0.003) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.255 (0.159) - 0.084 (0.005) - CMHE ($K=L$) 0.345 (0.104) 0.243 (0.042) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
+Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation.
+Best performance per column and simulation scenario are marked in bold, second best in italic.
+’-’ describes when the methodology diverges.
+Our proposed CSC method and its unadjusted variant best recover the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
+CSC outperforms all CMHE alternatives, the current state-of-the-art method.
+CMHE’s parametrisation and assumptions explain this difference.
+Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log hazards.
+Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data.
+In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes.
+Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods.
+These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
+CSC presents the best performance in identifying the underlying subgroup.
+Our proposed CSC method outperforms all approaches on the Rand-Index.
+In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are unaligned.
+KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index.
+Consequently, the identified clusters differ in their treatment responses, as shown by the large IAE.
+The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE.
+However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates.
+This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
+Treatment assignment correction improves subgroup identification in observational settings.
+The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation.
+The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context.
+However, CSC better recovers the different groups as shown by the Rand-Index.
+Finally, Appendices B.5.1 to B.5.4 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
+Real-world heterogeneity analysis in adjuvant radiotherapy responses (App. C).
+Using the Surveillance, Epidemiology, and End Results program(Seer) [^7] Available at https://seer.cancer.gov/, App.
+C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research.
+Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment.
+Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis.
+Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
 5 Conclusion
-This work fills the current gap in the literature of methodologies to identify subgroups of treatment responses in observational time-to-event data.
-While modelling the observed outcome, the approach jointly identifies patient subgroups with different treatment responses.
-Our experiments demonstrate the capacity of our proposed method to uncover groups of treatment effects in both synthetic and real-world datasets.
-However, causal modelling relies on empirically unverifiable assumptions.
-The invalidity of these assumptions is a lesser concern in our targeted setting of hypothesis generation, in comparison to individual estimation of treatment recommendations.
-Serving the purpose of a hypothesis-generating tool, we invite practitioners to further investigate the subgroups identified in observational studies through RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
+Understanding heterogeneity in treatment effect is essential for guiding clinical practice and informing the development of new treatments.
+Motivated principally by how real-world clinical guidelines depend on patient subgroups, this work introduces a neural network-based framework for identifying treatment-response subgroups in observational time-to-event data.
+By leveraging large-scale observational datasets and survival analysis with causal inference methods, our approach addresses a key gap in the literature: it moves beyond average or individualised treatment-effect estimation to reveal subgroups of treatment response.
+Our empirical results on both synthetic and real-world datasets demonstrate that the proposed method successfully characterises latent subgroups with divergent treatment effects.
+While we acknowledge that our model depends on empirically unverifiable assumptions as often in causal inference, our focus on subgroup detection for hypothesis generation lessens this concern.
+Indeed, such subgroup identification serves as a starting point for more targeted RCTs, where further validation can (i) confirm the heterogeneity in responses, (ii) explore alternative treatment strategies in less responsive subgroups, and (iii) shape evidence-based clinical guidelines.
+In doing so, this work contributes a flexible and practical tool for the design of future RCTs.
+By informing both the selection of patient populations most likely to benefit and avoiding treatments with limited or harmful effects, our framework offers a meaningful hypothesis-generating tool for clinicians, policymakers, and researchers striving to optimise patient care and resource allocation.
 Acknowledgements
-The authors acknowledge the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2).
+The authors acknowledge the partial support of the UKRI Medical Research Council (programme numbers MC_UU_00002/5 and MC_UU_00002/2 and theme number MC_UU_00040/02 – Precision Medicine).
 VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
@@
 A.1 Individualised Treatment Effect
-This section derives the individualised treatment effect expression introduced in equation 3.
+This section derives the individualised treatment effect expression introduced in Section 3.
 $\tau(t,x):=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)-\mathbbm{1}(T_{0}\geq t)\mid X=x)$
-$=\mathbb{E}(\mathbbm{1}(T_{1}\geq t)\mid X=x)-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
-$=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under assumption 4 and 5)
-A.2 Upper-Bound Simplification
-In Section 2, we claim that under the considered DAG presented in Figure 2,
-$\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$
-with $A$, being the treatment assignment. Under assumption 3, expressing this quality corresponds to:
-$q_{Z\mid X}^{A=a}:=q(Z\mid X,A=a)=q(Z\mid X)$
-with $q$, the density function. From this expression, $q_{Z\mid X}^{A=0}=q_{Z\mid X}^{A=1}$, which results in the distance between these distributions being null.
+$=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)\mid X=x)$
+$-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
+$=\mathbb{P}($ $T^{\prime}\geq t\mid A=1,X=x)$
+$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under assumption 1 and 3)
 Appendix B Synthetic analysis
@@
 We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups.
-The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to Nagpal et al.
-(2022a).
+The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to (Nagpal et al., 2022a).
 The following describes our generation process:
@@
 Note that we introduce $Y$ to present a covariates structure that is independent of the treatment responses.
+(a) $K=2$
+(b) $K=3$
+(c) $K=4$
+Figure 5: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross validated average estimated treatment effect. Lines in grey corresponds to the ground truth.
 Treatment response.
@@
 $\Gamma^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
-$T_{0}\mid Z,X,B^{0}_{z},\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
+$T_{0}\mid Z,X,B^{0}_{z},$ $\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$
+$\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
 $B^{1}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
 $\Gamma^{1}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
-$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{1}_{z}=(z,x,\beta_{z}^{1},\gamma_{z}^{1})$ $\sim\text{Gompertz}\left(w_{1}(\beta_{z}^{1},x),s_{1}(\gamma_{z}^{1},x)\right)$
-with $w_{0}$, $w_{1}$ two functions parametrising the Gompertz distributions’ shape as $w_{0}(\beta,x):=|\beta[0]|+(x[5:10]\cdot\beta[5:10])^{2}$, $w_{1}(\beta,x):=|\beta[0]|+(x[1:5]\cdot\beta[1:5])^{2}$, and the shift parameter parameterised as $s_{0}(\gamma,x):=|\gamma[0]|+|(x[1:5]\cdot\gamma[1:5])|$ and $s_{1}(\gamma,x):=|\gamma[0]|+|(x[5:10]\cdot\gamma[5:10])|$ where $v[i]$ described the $i^{th}$ element of the vector $v$.
+$T_{1}\mid Z,X,B^{1}_{z},$ $\Gamma^{1}_{z}=(z,x,\beta_{z}^{1},\gamma_{z}^{1})$
+$\sim\text{Gompertz}\left(w_{1}(\beta_{z}^{1},x),s_{1}(\gamma_{z}^{1},x)\right)$
+with $w_{0}$, $w_{1}$ two functions parametrising the Gompertz distributions’ shape as $w_{0}(\beta,x):=|\beta[0]|+(x[5:10]\cdot\beta[5:10])^{2}$, $w_{1}(\beta,x):=|\beta[0]|+(x[1:5]\cdot\beta[1:5])^{2}$, and the shift parameter parametrised as $s_{0}(\gamma,x):=|\gamma[0]|+|(x[1:5]\cdot\gamma[1:5])|$ and $s_{1}(\gamma,x):=|\gamma[0]|+|(x[5:10]\cdot\gamma[5:10])|$ where $v[i]$ described the $i^{th}$ element of the vector $v$.
 These functions aim to introduce non-linear responses with discrepancies between control and treatment regimes.
@@
 Treatment assignment.
-The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world applications.
+The non-randomisation of treatment is central to the problem of identifying treatment subgroups in real-world studies.
 Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment.
-We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as "Randomised" and "Observational".
-"Randomised" consists of a Bernoulli draw using the realisation of $P$.
-"Observational" reflects an assignment dependent upon the observed covariates.
-$A_{rand}\mid P=p$ $\sim\text{Bernoulli}(0.5)$
-$A_{obs}\mid P,X=(p,x)$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times 0.5)$
+We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as Randomised and Observational.
+Randomised consists of a Bernoulli draw using the realisation of $P$.
+Observational reflects an assignment dependent upon the observed covariates.
+$A_{rand}$ $\sim\text{Bernoulli}(0.5)$
+$A_{obs}\mid X=x$ $\sim\text{Bernoulli}(F_{\Phi(X)}(\Phi(x))\times 0.5)$
 with $F_{\Phi(X)}(\Phi(x))$ the cumulative distribution function that returns the probability that a realisation of a $\Phi(X)$ will take a smaller value than $\Phi(x)$.
-In our experiment, we chose $\Phi(x)=\sum x^{2}$.
+In our experiment, we chose $\Phi(x)=\sum x^{2}$ for a non-linear treatment assignment.
 Censoring.
@@
 with $w_{c}:=(x[5:10]\cdot\beta)^{2}$, the scale of the censoring Gompertz distribution.
-From these data, the goal is to model the treatment effect from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)*t_{0}+a*t_{1})$ and $D=\mathbbm{1}_{C>T}$.
+Our goal is to model the treatment effect and identify the underlying subgroup structure $Z$ from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)\times t_{0}+a\times t_{1})$ and $D=\mathbbm{1}_{C>T}$.
 B.2 Training and hyperparameter optimisation
@@
 We perform a 5-fold cross-validation for both Randomised and Observational simulations.
-For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2.
+For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search.
 All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
 Hyperparameter optimisation.
-We adopted a 100-iteration random grid-search over the following hyperparameters: network depth between 1 and 3 inner layers with 50 nodes, latent subgroup representation in $[10,25,50]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $100$ or $250$.
-B.3 Sensitivity
+We adopted a 500-iteration random grid-search over the following hyperparameters: network depth between 0 and 3 inner layers with 50 nodes, latent subgroup representation in $[10,25,50]$ and for training, a learning rate of $0.001$ or $0.0001$ with batch size of $100$ or $250$.
+The same set of parameters, when appropriate, was used for CMHE.
+Tree-based approach to estimate treatment effect had a depth limited to: 3, 5 or unconstrained, and the minimum sample split was chosen between 6, 12 or 24.
+B.3 Sensitivity to choice of K
 In Section 4, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$.
-This section explores how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters.
+While this outcome-guided selection of the number of clusters is a core strength of the proposed methodology in comparison to existing strategies, we explore how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters.
 This analysis examines the risk associated with misspecifying the number of clusters.
-Figure 5 illustrates the cross-validated subgroup treatment effects when trained with $K=2,3$ and $4$ clusters.
-Uncertainty increases with misspecified models due to the instability of the identified subgroups.
-(a) $K=2$
-(b) $K=3$ (as in the main text)
-(c) $K=4$
-Figure 5: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross validated average estimated treatment effect. Lines in grey corresponds to the ground truth.
-B.4 Alternative data generations
-In this section, we explore alternative data generations to demonstrate the robustness of the proposed strategy.
-B.4.1 Unequal cluster size
-Section B.1 presents a population equally distributed over the different clusters.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Obs.
+CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CSC Linear 0.819 (0.041) 0.047 (0.015) 0.031 (0.022) 0.022 (0.003)
+Table 2: Cross-validated performance (with standard deviation in parenthesis).
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.643 (0.067) 0.060 (0.016) 0.021 (0.007) 0.070 (0.013) CSC Unadjusted 0.643 (0.071) 0.064 (0.013) 0.018 (0.008) 0.060 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.788 (0.113) 0.040 (0.017) 0.041 (0.017) 0.086 (0.043) CSC Unadjusted 0.626 (0.170) 0.080 (0.015) 0.031 (0.009) 0.095 (0.049) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
+Table 3: Cross-validated performance (with standard deviation in parenthesis) when clusters are of different sizes.
+Figure 5 illustrates the cross-validated subgroup treatment effects when the model is trained with $K=2,3$ and $4$ clusters, despite the existence of 3 clusters.
+When using a number of clusters different than indicated by the elbow heuristic, the identified clusters are less stable as not capturing the true underlying treatment responses, as shown by the increased standard deviation.
+Uncertainty increases with misspecified models, providing an additional tool to select the appropriate number of cluster: the stability of the estimated clusters over cross-validation indicates a better alignment with underlying distribution.
+B.4 Linear Modelling
+While the flexibility of neural network allows flexibility of the survival distribution, one could consider more interpretable assignment $G$ and treatment $W$ components than the neural network approach proposed in the main text.
+This section explores the recovery of treatment effects and subgroups when using linear regressions, i.e.
+no hidden layers in a neural network, instead of the previous neural networks.
+Table 2 summarises the results presented in the main paper with the additional linear alternative in which both the treatment and assignment networks are replaced by a linear model, referred as CSC Linear.
+These results demonstrate that enforcing linear relation between covariates and the different quantities of interest can reduce the risk of overfitting in small datasets as shown by the improved Rand-Index.
+Note that under larger dataset or more complex relation between covariates, group membership and treatment, one should consider more flexible modelling, such as the proposed neural approach.
+B.5 Alternative data generations
+In this section, we explore alternative data-generative processes to demonstrate the robustness of the proposed strategy under different settings.
+B.5.1 Unequal cluster size
+Section B.1 presents an equally-distributed population over the different clusters.
 In medical settings, the underlying subgroups may differ in size.
-This section explores an alternative scenario in which the population is distributed over the three clusters: 62.5%, 25% and 12.5%.
+This section explores an alternative scenario in which the population is distributed over the three clusters as follows: 62.5%, 25% and 12.5%.
 Table 3 summarises the performance in this simulation, echoing the main text’s conclusions.
-The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects..
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.637 (0.068) 0.064 (0.013) 0.021 (0.009) 0.062 (0.010) CSC Unadjusted 0.651 (0.049) 0.055 (0.009) 0.017 (0.010) 0.067 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.780 (0.086) 0.064 (0.027) 0.040 (0.011) 0.069 (0.010) CSC Unadjusted 0.616 (0.146) 0.072 (0.020) 0.039 (0.028) 0.117 (0.033) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
-Table 2: Cross-validated performance (with standard deviation in parenthesis) when clusters are of different sizes.
-B.4.2 Number of points
+The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.310 (0.187) 0.236 (0.103) 0.110 (0.033) 0.149 (0.106) CSC Unadjusted 0.443 (0.285) 0.176 (0.098) 0.087 (0.046) 0.132 (0.112) CMHE ($K=3$) 0.113 (0.100) 0.296 (0.040) 0.137 (0.042) 0.188 (0.080) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.079 (0.119) 0.359 (0.068) 0.182 (0.089) 0.240 (0.058) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.764 (0.015) 0.027 (0.008) 0.011 (0.001) 0.047 (0.008) CSC Unadjusted 0.729 (0.022) 0.033 (0.008) 0.010 (0.003) 0.039 (0.009) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
+Table 4: Cross-validated performance (with standard deviation in parenthesis) with varying $N$ under observational treatment setting.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.830 (0.047) 0.051 (0.008) 0.023 (0.013) 0.020 (0.005) CSC Unadjusted 0.857 (0.033) 0.045 (0.006) 0.030 (0.015) 0.015 (0.003) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.134 (0.090) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.718 (0.198) 0.061 (0.050) 0.057 (0.048) 0.023 (0.005) CSC Unadjusted 0.804 (0.045) 0.037 (0.006) 0.023 (0.008) 0.025 (0.007) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.318 (0.140) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012)
+Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates under observational settings.
+B.5.2 Number of points
 Section B.1 describes a data generation with $3,000$ patients.
 This section presents results when there are $300$ and $30,000$ patients.
-A smaller number of patients may result in a larger impact of treatment non-randomisation but also impact the neural network capacity to identify underlying subgroups of treatment effect.
+A smaller number of patients may impact the neural network capacity to identify underlying subgroups of treatment effect.
 A larger number makes non-randomisation less of a concern.
@@
 However, throughout the different settings, the proposed CSC presents the best Rand-Index indicating a good recovery of the underlying structure.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.504 (0.139) 0.131 (0.090) 0.085 (0.033) 0.091 (0.062) CSC Unadjusted 0.449 (0.106) 0.202 (0.049) 0.162 (0.055) 0.125 (0.088) CMHE ($K=3$) 0.126 (0.084) 0.340 (0.050) 0.135 (0.043) 0.166 (0.050) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.087 (0.118) 0.284 (0.056) 0.639 (0.939) 0.747 (1.014) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.774 (0.015) 0.018 (0.010) 0.012 (0.003) 0.030 (0.009) CSC Unadjusted 0.737 (0.022) 0.033 (0.006) 0.010 (0.002) 0.039 (0.008) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
-Table 3: Cross-validated performance (with standard deviation in parenthesis) with varying $N$ under observational treatment setting.
-B.4.3 Number of clusters
-Section B.1 describes a data generation with $K=3$ clusters.
-This section explores the behaviour of the different methodologies with an increased number of clusters $K=5$.
-Following the same data generation, we introduces two additional clusters with centres [-3, 3] and [4, 4].
-Table 4 presents the different models’ performance in this setting, echoing the same results as the main text.
-The proposed method best recovers the underlying clustering structure as shown by the Rand-Index.
-CMHE suffers from its inherent association between treatment and survival.
-The step-wise approach is unable to identify the clusters of interest if the data presents a structure that is not fully aligned with the outcome of interest.
-This scenario further validates our findings, demonstrating the methodology’s capacity to generalise beyond the scenario presented in the main text.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ Randomised CSC 0.475 (0.038) 0.073 (0.030) 0.059 (0.023) 0.041 (0.015) 0.099 (0.024) 0.017 (0.003) CSC Unadjusted 0.459 (0.037) 0.095 (0.055) 0.054 (0.027) 0.046 (0.011) 0.076 (0.020) 0.019 (0.005) CMHE ($K=5$) 0.306 (0.016) 0.091 (0.036) 0.104 (0.012) 0.015 (0.002) 0.093 (0.011) 0.232 (0.013) CMHE ($L=5$) 0.256 (0.018) 0.135 (0.020) 0.112 (0.008) 0.020 (0.003) 0.221 (0.009) 0.349 (0.006) CMHE ($K=L$) 0.372 (0.018) 0.101 (0.023) 0.181 (0.013) 0.072 (0.003) 0.119 (0.010) 0.323 (0.010) KMeans + TE 0.156 (0.012) 0.075 (0.009) 0.184 (0.016) 0.072 (0.012) 0.147 (0.014) 0.039 (0.010) Virtual Twins 0.284 (0.081) 0.034 (0.029) 0.066 (0.026) 0.029 (0.012) 0.082 (0.023) 0.036 (0.010) Observational CSC 0.446 (0.016) 0.119 (0.072) 0.087 (0.018) 0.026 (0.008) 0.097 (0.023) 0.027 (0.016) CSC Unadjusted 0.444 (0.048) 0.077 (0.046) 0.065 (0.026) 0.045 (0.020) 0.118 (0.026) 0.034 (0.022) CMHE ($K=5$) 0.303 (0.010) 0.095 (0.030) 0.112 (0.017) 0.018 (0.005) 0.094 (0.013) 0.229 (0.010) CMHE ($L=5$) 0.217 (0.026) 1.341 (2.717) 0.119 (0.014) 0.022 (0.005) 1.432 (2.720) 1.561 (2.709) CMHE ($K=L$) 0.371 (0.029) 0.105 (0.009) 0.208 (0.043) 0.076 (0.006) 0.186 (0.025) 0.325 (0.004) KMeans + TE 0.162 (0.011) 0.058 (0.005) 0.210 (0.031) 0.084 (0.018) 0.122 (0.032) 0.032 (0.006) Virtual Twins 0.320 (0.084) 0.062 (0.038) 0.026 (0.008) 0.044 (0.021) 0.077 (0.040) 0.038 (0.018)
-Table 4: Cross-validated performance (with standard deviation in parenthesis) when $K=5$.
-B.4.4 Impact of treatment rate
+B.5.3 Impact of treatment rate
 Section B.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.559 (0.114) 0.024 (0.009) 0.030 (0.017) 0.030 (0.006) CSC Unadjusted 0.481 (0.069) 0.026 (0.005) 0.041 (0.026) 0.032 (0.009) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.254 (0.138) 0.074 (0.027) 0.052 (0.021) 0.149 (0.018) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.584 (0.151) 0.015 (0.007) 0.028 (0.024) 0.035 (0.003) CSC Unadjusted 0.556 (0.249) 0.023 (0.008) 0.051 (0.025) 0.040 (0.006) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.293 (0.125) 0.063 (0.012) 0.051 (0.017) 0.140 (0.011) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
+Table 6: Cross-validated performance (with standard deviation in parenthesis) when all covariates clusters presents different treatment responses.
+This setting aligns with the assumptions made by the step-wise approaches.
 Table 5 presents the performance under these different treatment rates in an observational setting.
 These results highlight CSC’s capacity to identify the underlying subgroups with the highest Rand-Index and one of the best cluster treatment effect recovery in these settings.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.790 (0.036) 0.036 (0.009) 0.034 (0.009) 0.018 (0.006) CSC Unadjusted 0.849 (0.033) 0.040 (0.005) 0.019 (0.005) 0.014 (0.004) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.149 (0.080) 0.188 (0.004) - 0.154 (0.005) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.808 (0.043) 0.035 (0.012) 0.027 (0.006) 0.019 (0.007) CSC Unadjusted 0.687 (0.048) 0.045 (0.011) 0.035 (0.016) 0.028 (0.006) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.010) 0.077 (0.005) 0.069 (0.004) CMHE ($L=3$) 0.318 (0.140) - - 0.132 (0.008) CMHE ($K=L$) 0.483 (0.104) 0.208 (0.013) - - KMeans + TE 0.002 (0.003) 0.195 (0.015) 0.079 (0.012) 0.148 (0.011) Virtual Twins 0.570 (0.013) 0.030 (0.005) 0.024 (0.013) 0.022 (0.005)
-Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates.
-B.4.5 Aligned covariates and treatment effect structure
+B.5.4 Aligned covariates and treatment effect structure
 Section B.1 presents a data generation with the last 8 covariates having a clustered structures independent of the treatment response.
@@
 $X_{[3-10]}\sim\text{MVN}(0,I^{8})$
-Table 6 summarises the performance in this setting, evidencing an improvement of the Kmeans+TE approach.
+Table 6 summarises the performance in this setting, evidencing an improvement of the Kmeans+TE approach as this model makes this assumption.
 When covariates’ clusters are aligned with the outcome of interest, this method recovers the clustering structure well, as shown by the best Rand-Index.
-Non-randomisation of treatment, however, worsens the treatment estimates due to the absence of correction in the estimation of treatment responses.
 Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses.
-This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across settings.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.487 (0.052) 0.027 (0.013) 0.047 (0.020) 0.030 (0.009) CSC Unadjusted 0.522 (0.041) 0.025 (0.004) 0.027 (0.018) 0.034 (0.013) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.289 (0.078) 0.063 (0.009) 0.050 (0.019) 0.146 (0.020) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.666 (0.146) 0.016 (0.007) 0.029 (0.013) 0.033 (0.010) CSC Unadjusted 0.613 (0.207) 0.024 (0.016) 0.039 (0.015) 0.030 (0.003) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.294 (0.123) 0.057 (0.005) 0.061 (0.007) 0.144 (0.007) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
-Table 6: Cross-validated performance (with standard deviation in parenthesis) when all covariates clusters presents different treatment responses.
-This setting aligns with the assumptions made by the step-wise approaches.
-B.4.6 Linear treatment responses
-As an alternative setting, we propose a setting with a linear treatment response, as assumed by CMHE Nagpal et al.
-(2023).
-To this end, we followed a similar process as described in Section B.1 at the exception of the treatment response.
-We draw the response under no treatment from a Gompertz, and the one under treatment with the same parameters shifted by the treatment effect.
-Formally, using the same notation
-$B^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
-$\Gamma^{0}_{z}\mid Z=z$ $\sim\text{MVN}(0,I^{10})$
-$T_{0}\mid Z,X,B^{0}_{z},\Gamma^{0}_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x),s_{0}(\gamma_{z}^{0},x)\right)$
-$T_{1}\mid Z,X,B^{1}_{z},\Gamma^{0}_{z},A_{z}=(z,x,\beta_{z}^{0},\gamma_{z}^{0},a_{z})$ $\sim\text{Gompertz}\left(w_{0}(\beta_{z}^{0},x)\times a_{z},s_{0}(\gamma_{z}^{0},x)\right)$
-with $a_{z}$ the treatment effect for cluster $z$. In this experiments, we choose $a_{0},a_{1},a_{2}=(0.5,1,1.5)$.
-Table 7 summarises performances in this scenario.
-While demonstrating a significant improvement of the CMHE alternatives, CSC presents competing performance on subgroup identification and treatment effect estimates.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.389 (0.065) 0.033 (0.015) 0.058 (0.035) 0.014 (0.006) CSC Unadjusted 0.411 (0.125) 0.029 (0.017) 0.062 (0.021) 0.017 (0.015) CMHE ($K=3$) 0.304 (0.033) 0.040 (0.003) 0.050 (0.003) 0.030 (0.003) CMHE ($L=3$) 0.211 (0.109) 0.051 (0.003) 0.106 (0.006) 0.028 (0.008) CMHE ($K=L$) 0.384 (0.060) 0.411 (0.797) - 0.435 (0.650) KMeans + TE 0.101 (0.139) 0.080 (0.011) 0.064 (0.028) 0.038 (0.023) Virtual Twins 0.265 (0.085) 0.025 (0.006) 0.018 (0.004) 0.064 (0.013) Observational CSC 0.311 (0.043) 0.029 (0.003) 0.074 (0.028) 0.027 (0.016) CSC Unadjusted 0.275 (0.168) 0.032 (0.005) 0.044 (0.044) 0.026 (0.014) CMHE ($K=3$) 0.360 (0.068) 0.038 (0.002) 0.050 (0.002) 0.035 (0.009) CMHE ($L=3$) 0.285 (0.111) 0.055 (0.006) 0.094 (0.007) 0.017 (0.009) CMHE ($K=L$) 0.382 (0.058) 0.052 (0.011) 0.128 (0.013) 0.028 (0.020) KMeans + TE 0.115 (0.154) 0.068 (0.020) 0.076 (0.031) 0.031 (0.008) Virtual Twins 0.186 (0.117) 0.048 (0.027) 0.041 (0.010) 0.039 (0.012)
-Table 7: Cross-validated performance (with standard deviation in parenthesis) when considering a linear treatment response. This setting aligns with the assumption made by CMHE.
-Appendix C Real-world case Study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
+This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across different and even adverserial settings.
+Appendix C Case study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
 To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017.
-Following Lee et al.
-(2018); Danks & Yau (2022); Jeanselme et al.
-(2023), we select women who died from the condition or from cardiovascular diseases.
+Following (Lee et al., 2018; Danks & Yau, 2022; Jeanselme et al., 2023), we select women who died from the condition or from cardiovascular diseases.
 From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes.
 To this end, we subselect patients with recorded treatment and who received chemotherapy.
-These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  Danks & Yau (2022) for further description).
+These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  (Danks & Yau, 2022) for further description).
 Selection of $K$.
@@
 In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy.
-This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist Lazzari et al.
-(2023), making this setting more likely to meet the positivity assumption (assumption 6), necessary to study causality in observational data.
-(a) Neural Survival Treatment
+This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist (Lazzari et al., 2023), making this setting more likely to meet the positivity assumption (Assumption 4), necessary to study causality in observational data.
+(a) Causal Survival Clustering
 (b) Cox Mixture of Heterogeneous Effects
 Figure 7: Averaged treatment effect subgroups across 5-fold cross-validation observed in the Seer dataset with the shaded areas representing 95% CI.
+Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
+RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.01 (0.00) 94.6% 55.6% 1.2 (5.90) 17.5% 46.5% Subgroup 1 0.84 (0.11) 5.4% 45.1% 20.5 (15.56) 23.4% 50.4%
+Table 7: Causal Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
 Uncovering treatment response.
-Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE, with the number of subgroups $K$ selected through hyperparameter tuning.
+Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE.
+For this latter, the number of subgroups $K$ is selected through hyperparameter tuning, leading to $K=1$ in every fold.
 As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups.
@@
 Figure 8 displays the 10 covariates most indicative of the different treatment response subgroups.
-Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
-Table 8 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) Royston & Parmar (2013).
+Table 7 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) (Royston & Parmar, 2013).
 Both methodologies identify a population with limited treatment response.
 However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
-RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.00 (0.00) 93.3% 55.5% 1.18 (5.74) 17.4% 46.6% Subgroup 1 0.62 (0.16) 6.7% 47.5% 17.75 (16.08) 23.4% 48.6%
-Table 8: Causal Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
 The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy.

### v5

@@
 Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials.
-Existing approaches for treatment effect estimation primarily rely on Randomised Controlled Trials (RCTs), which are often limited by insufficient power, multiple comparisons, and unbalanced covariates.
-In addition, RCTs tend to feature more homogeneous patient groups, making them less relevant for uncovering subgroups in the population encountered in real-world clinical practice.
+Existing approaches for treatment effect estimation primarily rely on Randomised Controlled Trials (RCTs), which tend to feature more homogeneous patient groups, making them less relevant for uncovering subgroups in the population encountered in real-world clinical practice.
 Subgroup analyses established for RCTs suffer from significant statistical biases when applied to observational studies, which benefit from larger and more representative populations.
@@
 In experiments, our approach significantly outperforms the current state-of-the-art method for subgroup analysis in both randomised and observational treatment regimes.
-Machine Learning, ICML
+keywords:
+Treatment effect, subgrouping, observational data
+Data and Code Availability
+Experiments are performed on synthetic data and a publicly available dataset from the Surveillance, Epidemiology, and End Results Program [^1] Available at https://seer.cancer.gov/.
+The code to reproduce the proposed model and the presented results is available on GitHub [^2] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
+Institutional Review Board (IRB)
+This research does not require IRB approval as it relies on a publicly available dataset from previously approved studies.
 1 Introduction
@@
 By randomly assigning patients to control or treated arms, RCTs eliminate potential confounders when assessing a treatment’s impact on outcomes.
-However, RCTs are often costly, time-consuming, and restricted to specific patient cohorts, which are likely not to reflect the full spectrum of patients seen in clinical practice (Hernán & Robins, 2016, 2010; Cole & Hernán, 2008).
-Trial findings may thus not generalise to real-world settings, where treatment choice depends on clinical judgment and patient preferences, among other variables that may remain unobserved.
+However, RCTs are often costly, time-consuming, and restricted to specific patient cohorts, which are likely not to reflect the full spectrum of patients seen in clinical practice (Hernán and Robins, 2016, 2010; Cole and Hernán, 2008).
+Trial findings may thus not generalise to real-world settings.
 Figure 1: Subgroup treatment effect discovery in time-to-event observational data.
 Our work aims to identify subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials.
-Our method simultaneously models the treatment effect and identifies subgroup while addressing treatment non-randomisation and censoring.
+Our method simultaneously models the treatment effect and identifies subgroups while addressing treatment non-randomisation and censoring.
 To address these limitations, researchers increasingly leverage large-scale observational datasets which capture diverse, real-world populations.
-Although observational data offer the potential to identify subgroups that RCTs might miss, they also pose challenges due to non-random treatment assignments and confounding (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018).
+Although observational data offer the potential to identify subgroups that RCTs might miss, they also pose challenges due to non-random treatment assignments and confounding (Benson and Hartz, 2000; Hernán and Robins, 2010; Hernán, 2018).
 A robust methodological literature in causal inference has emerged to tackle these challenges.
-Techniques such as inverse probability weighting (Cole & Hernán, 2008), marginal structural models (Robins et al., 2000), and doubly robust estimators (Bang & Robins, 2005) allow for more accurate causal effect estimation under appropriate assumptions.
+Techniques such as inverse probability weighting (Cole and Hernán, 2008), marginal structural models (Robins et al., 2000), and doubly robust estimators (Bang and Robins, 2005) allow for more accurate causal effect estimation under appropriate assumptions.
 Prior works in machine learning (ML) have extended these ideas to estimate treatment effects from observational data (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017).
-However, most methods focus on (i) average treatment effects at the population level or (ii) individualised treatment effects.
-These approaches overlook the utility of identifying treatment-effect subgroups —groups with consistently distinct responses to a given intervention.
-Detecting such subgroups is crucial in translating treatment effect estimates into actionable insights for clinical decision-making and resource allocation.
+However, most methods focus on (i) average treatment effects (ATE), which measure the response at the population level, (ii) conditional average treatment effect (CATE), which reflects average response for an individual given its covariates, or (iii) individualised treatment effects (ITE), which aims to estimate the unobservable counterfactual effect.
+These approaches overlook the utility of identifying treatment-effect subgroups —groups with distinct responses to a given intervention.
+Identifying such subgroups is crucial in translating treatment effect estimates into actionable insights for clinical decision-making and resource allocation.
 We address this gap by introducing a novel framework that uncovers patient subgroups with distinct treatment responses using observational time-to-event data.
-Extending the rich tradition of outcome-guided modelling (Foster et al., 2011) to a flexible, neural network–based architecture.
+Extending the rich tradition of outcome-guided modelling (Foster et al., 2011) to a flexible, neural network–based architecture, our proposed approach addresses the limitations of existing methodology for subgroup discovery (see App.
+A for a full review).
 Crucially, our approach does not require explicit parametrisation of time-to-event distributions or treatment effects; it instead leverages a mixture of monotonic neural networks with correction for non-random treatment assignment.
 In summary, our contributions are:
+  • Novel formalisation. In Sec. 2, we formalise the problem of treatment subgroup discovery.
   • Neural network-based treatment subgroup identification.
-Our work introduces the first neural network architecture to simultaneously estimate and identify subgroups of treatment effects in observational settings, accommodating both censoring and non-random treatment assignments without parameterising the treatment effects.
-  • Novel formalisation.
-In section 2, we review existing approaches to identify subgroups of treatment effects.
-Section 3 formalises the problem of subgroup discovery and our strategy for minimising biases arising from non-random treatment assignment.
-  • Extensive application on survival time-to-event data.
-In section 3, we introduce the proposed monotonic neural network implementation, and evaluate its performance on two synthetic datasets in Section 4, an extensive set of alternative settings in App.
-B and a real-world example described in App.
-C, benchmarking against several established approaches.
-2 Related work
+We introduce, in Sec.
+2, a neural network architecture to jointly estimate and identify subgroups of treatment effects in observational settings.
+  • Extensive application on time-to-event data.
+We evaluate its performance on synthetic data in Sec.
+3, with extensive sensitivity analysis in App.
+C, and a real-world example described in Sec.
+4, benchmarking against several established approaches.
+2 Method
+2.1 Problem setup
+Our goal is to uncover a pre-specified $K\in\mathbb{N}$ number of subgroups [^3] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Sec.
+3.
+with distinct treatment responses, guided by the observed times of occurrence of an outcome of interest.
+To this end, our model assigns patients to latent subgroups based on their covariates at the time of treatment decision.
+For each subgroup, we estimate the probability of not observing the event of interest over time under treatment and under control regimes.
+These two distributions are known as survival functions, and their difference corresponds to the estimated treatment effect for a given group.
+Formally, consider the random variables associated with observed covariates $X$, the assigned binary treatment $A$, and the observed event time $T^{\prime}$.
+Following the potential outcomes formulation, $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under consistency (Asmp.
+1) where $T^{\prime}_{1}$ is the potential event time under treatment and $T^{\prime}_{0}$, under the control regime.
+Assumption 1 (Consistency)
+A patient’s observed event time is the potential event time associated with the observed treatment.
+Formally, this means $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the control and treatment regimes, respectively.
+Central to our problem is the latent, unobserved, subgroup membership $Z$.
+As formalised in Asmp.
+2, we assume that the variable $Z$ mitigates the dependence between the covariates $X$ and the observed event time $T^{\prime}$.
+Assumption 2 (Mixture mitigation)
+The event time $T^{\prime}$ is determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
+In this setting, our interest lies in estimating the Subgroup Average Treatment Effect (SATE) for group $k$:
+$\tau_{k}(t)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,Z=k)$
+$-\mathbb{P}(T^{\prime}\geq t\mid A=0,Z=k)$ (1)
+This focus differs from the literature interest in estimating CATE, connected to the previous quantity as follows (see App. B for derivation):
+$\tau(t,x)$ $=\sum_{k}\tau_{k}(t)\mathbb{P}(Z=k\mid X=x)$
+Estimating the previous quantities requires the accurate modelling of the survival distributions under the two treatment regimes.
+The central challenge in this estimation is that the counterfactual survival outcome is unobserved: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
+The absence of treatment randomisation in observational studies, e.g.
+clinicians might recommend more aggressive treatment for more severe conditions, results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021).
+This prohibits the estimation of the survival functions through the maximisation of the likelihood of the observed outcomes.
+Under the common Asmp.
+3 and 4 (Hernán and Robins, 2010; Robins, 1986), methods exist to account for the difference between the treated and non-treated populations.
+One such method adopted in this work is the Inverse Propensity Weighting (IPW), which consists of weighting each observation by the inverse probability of receiving treatment to estimate the overall likelihood.
+Assumption 3 (Conditional ignorability)
+The potential event times are independent of the treatment given the observed covariates, i.e.
+$A\perp\!\!\!\perp(T^{\prime}_{0},T^{\prime}_{1})\mid X$.
+Equivalently, no unobserved confounders impact both treatment and event time.
+Assumption 4 (Overlap / Positivity)
+Each patient has a non-zero probability of receiving the treatment, i.e. $\mathbb{P}(A\mid X)\in(0,1)$ where $(0,1)$ is the open interval, resulting in a non-deterministic treatment assignment.
+A final challenge exists in time-to-event data: censoring.
+Patients may not observe the outcome of interest during the study period.
+Formally, instead of observing $T^{\prime}$, one observes an indicator identifying whether the event of interest was observed $D$, and the associated observed time of event $T$, with $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is the random variable of the (right)-censoring time.
+When a patient is censored, $T=C$ and $D=0$; otherwise, the event under treatment regime $A$ is observed and $T=T^{\prime}_{A}$ and $D=1$.
+As ignoring the censoring process biases the likelihood, we assume, as commonly done in survival analysis, that this process is uninformative, formalised as:
+Assumption 5 (Non-informative censoring)
+The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$, given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
+Under the previous assumptions, common in the causal and survival literature, we can maximise the overall likelihood by optimising for the following weighted factual log-likelihood $l_{F}$:
+$l_{F}$ $=\sum_{i,d_{i}=1}w_{i}\log\left(-\frac{\partial S(t\mid x_{i},a_{i})}{\partial t}\bigg|_{t=t_{i}}\right)$
+$+\sum_{i,d_{i}=0}w_{i}\log S(t_{i}\mid x_{i},a_{i})$ (2)
+where $i$ is the patient index, $(x_{i},t_{i},a_{i},d_{i})$ are realisations of the associated variables $X,T,A$ and $D$, and $w_{i}$ is the inverse propensity weighting correction for patient $i$.
+Under Asmp.
+2,
+$S(t$ $\mid x,a):=\mathbb{P}(T^{\prime}\geq t\mid x,a)$
+$=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid x)\mathbb{P}(T^{\prime}\geq t\mid a,k)$ (3)
+Figure 2 summarises all variables and dependencies assumed in the studied problem with a directed acyclic graph.
+Given the observed $X$, AA, $T$, and $D$, our aim is to estimate the latent structure $Z$ and the associated survival distributions under the two potential treatment regimes $T^{\prime}_{0}$ and $T^{\prime}_{1}$.
+Figure 2: Graphical representation between covariates ($X$), treatment (AA) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, AA, $T$ and $D$ are observed.
+2.2 Estimating the quantities of interest
+The previous section discussed the quantities one must estimate —here parametrised by neural networks [^4] When interpretability is a key concern, one can consider a linear model for both subgroup assignment and IPW components, as explored in App.
+C.4.— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the distributions characterising survival under the two treatment regimes $\mathbb{P}(T^{\prime}\geq t\mid A,Z=k)$, and the IPW weights $w$.
+Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
+Subgroup assignment.
+A multi-layer perceptron $G$ with a final Softmax layer assigns a patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
+$G(x):=[\mathbb{P}(Z=k\mid x)]_{k=1}^{K}$
+Survival distributions.
+Each subgroup $k$ is represented by a parameter vector $u_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation learnt through backpropagation.
+The vector $u_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic outcomes as in Jeanselme et al.
+(2022, 2023) and a final SoftPlus layer to ensure positivity.
+We train this neural network to model survival through the following transformation — ensuring that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
+$\mathbb{P}(T^{\prime}\geq t\mid A=a,Z=k):=\exp(-t\cdot M(u_{k},t)[a])$
+Inverse propensity weighting.
+As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity of receiving treatment estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as:
+$W(x):=\mathbb{P}(A=1\mid x)$
+From this estimated probability, we further truncate the propensity score (Austin and Stuart, 2015) to avoid extreme values for the estimated weights, which would result in unstable estimates of the treatment effect.
+In this context, the weights $w_{i}$ are defined as:
+$\forall i,w_{i}^{\prime}:$ $=a_{i}W(x_{i})+(1-a_{i})(1-W(x_{i}))$
+$w_{i}:$ $=\max(0.05,\min(w_{i}^{\prime},0.95))$
+Figure 3: Causal Survival Clustering (CSC) architecture.
+Latent parameter $u_{k}$ characterising the subgroup $k$ is inputted into the monotonic network $M$ to estimate the survival under both treatment regimes.
+$G$ assigns the probability of belonging to each subgroup given the patient’s covariate(s) $x$.
+The network $W$ estimates the treatment propensity used to account for the treatment assignment bias.
+2.3 Training procedure
+The model $W$ is trained to predict the binary treatment assignment probability by minimising the cross-entropy of receiving treatment.
+Then, all other components are trained by maximising the weighted log-likelihood introduced in Equation (2).
+The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood, as a forward path estimates survival and automatic differentiation provides the derivative necessary for computing the log-likelihood (Jeanselme et al., 2022, 2023; Rindt et al., 2022).
+3 Experimental analysis
+As counterfactuals are unknown in observational data, we adopt the common practice of a synthetic dataset in which the underlying survival distributions and group structure are known.
+As a real-world case study, App.
+4 accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer.
+Our code to produce the synthetic dataset, the model, and all experimental results is available on Github [^5] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
+3.1 Data generation
+We generate a population of $3,000$ equally divided into $K=3$ subgroups.
+We draw 10 covariates from normal distributions with different means and survival distributions for each treatment regime following Gompertz distributions (Pollard and Valkovics, 1992) parametrised by group membership and individual covariates.
+Note that this setting breaks Asmp.
+2 as survival distributions under treatment or control regimes are different functions of the covariates.
+This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis, which often assume a linear treatment response.
+Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational).
+Finally, non-informative censoring times are drawn following a Gompertz distribution.
+Details of the generative process for our synthetic dataset are deferred to App.
+C.1.
+Further, App.
+C.5 explores alternative datasets to demonstrate the robustness of the proposed methodology.
+3.2 Empirical settings
+Benchmark methods.
+We compare the proposed approach (CSC) against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE Nagpal et al.
+(2022a) [^6] Implemented in the Auton-Survival library (Nagpal et al., 2022b)), which uncovers treatment effect and baseline survival latent groups.
+This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted.
+The central differences with our proposed approach are that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response.
+This separation between survival ($M$) and treatment response ($L$) enhances the model’s flexibility at the expense of interpretability, as the total number of groups grows exponentially ($L\times M$), and subgroups of treatment effect become independent of survival.
+Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses.
+By contrast, our approach identifies treatment subgroups while considering survival without constraining the treatment response.
+We argue that these are key strengths of our method, as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, compared to a group with the same treatment responses but with better survival odds.
+Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
+For a fair comparison, we present three alternatives of CMHE: one with fixed $M=3$ survival subgroups and $L=1$ treatment subgroups, one with $L=3$ treatment and $M=1$ survival subgroups, and one with $M=L=2$, which allows for a total of 4 subgroups.
+We consider these alternatives to obtain a total number of clusters close to the underlying number of subgroups $K$.
+Crucially, CMHE assumes proportional hazards for each subgroup and does not account for the treatment non-randomisation.
+As an ablation, we compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and therefore is not robust to non-random treatment assignment.
+Finally, we compare against two step-wise approaches in which clustering and treatment effect are trained separately.
+First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in Nagpal et al.
+(2022a), referred to as KMeans + TE in the following.
+Specifically, we use a KMeans (Hartigan and Wong, 1979) to cluster the data, and we compute the difference between Kaplan-Meier (Kaplan and Meier, 1958) estimates between control and treated patients stratified by clusters.
+Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them.
+As proposed in the literature, we use a survival tree to estimate response under each treatment regime and then use a KMeans (MacQueen et al., 1967) on the estimated treatment effects, computed as the difference between survival estimates.
+For details on training and hyperparameter optimisation, we refer to App.
+C.2.
+Evaluation.
+In the synthetic experiments, the subgroup structure is known.
+We measure the adjusted [^7] Random patient assignment results in an adjusted Rand-Index of 0.
+Rand-Index (Rand, 1971), which quantifies how the estimated subgroup assignment aligns with the generative group structure.
+Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well the methods recover each subgroup’s treatment effect
+$\text{IAE}_{k}(t)=\int_{0}^{t}\left|\hat{\tau}_{\hat{k}}(s)-\tau_{k}(s)\right|ds,$
+where $\hat{\tau}_{\hat{k}}$ is the estimated treatment effect for subgroup $\hat{k}=\operatorname*{arg\,max}_{l}\mathbb{E}_{x\in k}(Z=l\mid x)$, i.e.
+the most likely assigned cluster for patients in the underlying $k$ cluster, and $\tau_{k}$ is the ground truth.
+3.3 Treatment effect recovery
+Recovering the underlying number of subgroups.
+For all methodologies, we must choose the number of subgroups $K$ a priori.
+An important question is, therefore, whether we can identify the underlying number of subgroups in a principled way using the proposed CSC.
+Figure 4 presents the average negative log-likelihood obtained by cross-validation for models with different numbers of subgroups $K$.
+Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the ”Observational” treatment assignment with the shaded area representing 95% CI.
+The log-likelihood presents an elbow around the underlying number of subgroups.
+The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability, here considering the log-likelihood.
+Using this heuristic, the optimal choice for $K$ is $3$, which aligns with the generative process.
+App.
+C.3 shows that when the number of clusters is misspecified, the estimated treatment effects are unstable; the variance associated with estimated treatment effects is hence an additional tool for informing the choice of $K$.
+These data-driven approaches to select the number of subgroups $K$ are a crucial strength of our method, compared to classical two-stage analyses, which separate clustering from treatment effect estimates.
+In these methods, survival outcomes cannot directly guide the choice of $K$, whereas it is directly reflected in CSC’s performance.
+Discovering subgroups.
+Table 1 summarises the performance of the different methodologies under the two studied scenarios.
+Recall that $L$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $M$ describes the number of survival clusters.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.807 (0.055) 0.037 (0.008) 0.023 (0.008) 0.015 (0.003) CSC Unadjusted 0.802 (0.050) 0.039 (0.011) 0.021 (0.009) 0.014 (0.003) CMHE ($L=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($M=3$) 0.255 (0.159) - 0.084 (0.005) - CMHE ($M=L$) 0.345 (0.104) 0.243 (0.042) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CMHE ($L=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($M=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($M=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
+Table 1: 5-fold cross-validated performance averaged (with standard deviation in parentheses) under the Randomised and Observational treatment simulation.
+Best performance per column and simulation scenario is marked in bold, second best in italic.
+’-’ describes when the methodology diverges.
+Our proposed CSC method best recovers the underlying treatment responses in the observational setting.
+CSC outperforms all CMHE alternatives, the current state-of-the-art method.
+CMHE’s parametrisation and assumptions explain this difference.
+Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log hazards.
+Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data.
+In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes.
+Increasing the number of subgroups, as shown in $M=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods.
+These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes.
+CSC presents the best performance in identifying the underlying subgroup.
+Our proposed CSC method outperforms all approaches on the Rand-Index.
+In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are unaligned.
+KMeans+TE identifies subgroups that are independent of treatment subgroups, as shown by the Rand-Index and the large IAE.
+The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE.
+However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates.
+This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
+Treatment assignment correction improves subgroup identification in observational settings.
+The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation.
+The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context.
+However, CSC better recovers the different groups as shown by the Rand-Index in the Observational setting.
+4 Case study: Analysis of adjuvant radiotherapy responses
+To illustrate how practitioners can use the proposed methodology to identify subgroups of treatment response in real-world observational data, we investigate how breast cancer patients respond differently to adjuvant radiotherapy using data from the Surveillance, Epidemiology, and End Results program (Seer).
+Following Lee et al.
+(2018); Danks and Yau (2022); Jeanselme et al.
+(2023), our analysis focuses on women who died from the condition or from cardiovascular diseases.
+To study the impact of adjuvant radiotherapy on survival outcomes after chemotherapy, we subselect patients with recorded chemotherapy.
+These criteria led to the selection of 239,855 women with 22 covariates measured at diagnosis, such as diagnosis year, grades, ethnicity, laterality, tumour size, and type (see Danks and Yau (2022) for further description).
+RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.01 (0.00) 94.6% 55.6% 1.2 (5.90) 17.5% 46.5% Subgroup 1 0.84 (0.11) 5.4% 45.1% 20.5 (15.56) 23.4% 50.4%
+Table 2: Causal Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
+Step 1: Select $K$.
+From this population, our aim is to identify heterogeneous responses to adjuvant treatment.
+The first challenge is selecting the number of groups to use ($K$).
+We advise following medical actionability and considering the change in treatment effects and size of the subgroups when increasing this parameter.
+In the absence of experts’ insights, one may rely on the previously described elbow rule heuristic over $l_{F}$.
+Using Figure 7 in the Appendix, the negative log-likelihood presents an elbow for $K=2$.
+Step 2: Identify subgroups.
+Once $K$ is selected, one may use CSC to identify groups of patients who may benefit from adjuvant radiotherapy.
+This problem is central to patients’ management, as no evidence-based guidelines for adjuvant therapy exist Lazzari et al.
+(2023), making this setting more likely to meet the positivity assumption (Asmp.
+4), which is necessary to study causality in observational data.
+Figure 5: Averaged treatment effect subgroups across 5-fold cross-validation observed in the Seer dataset with the shaded areas representing 95% CI.
+Figure 5 presents the identified treatment effect subgroups, and Table 2 summarises the mean covariates across the identified subgroups and the life expectancy gain when using adjuvant radiation, measured through the Restricted Mean Survival Time (RMST) (Royston and Parmar, 2013).
+The proposed methodology identifies two groups: one with limited treatment response and one characterised by larger HER2 prevalence and higher distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
+See App.
+D for the analysis of the treatment subgroups identified using the other baselines.
+Step 3: Validate.
+The proposed analysis identifies a group that benefits from adjuvant radiotherapy.
+However, our methodology remains a hypothesis-generating tool.
+An important next step would be to develop clinical trials to validate the identified subgroups, particularly due to potential confounding factors such as hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options.
+The quality of available covariates limits this analysis and serves only as an example for medical practitioners to identify subgroups of treatment responses from observational data.
+5 Conclusion
+Understanding heterogeneity in treatment effect is essential for guiding clinical practice and informing the development of new treatments.
+Motivated primarily by the reliance of clinical guidelines on patient subgroups, this work introduces a neural network-based framework for identifying treatment-response subgroups in observational time-to-event data.
+By leveraging large-scale observational datasets and survival analysis with causal inference methods, our approach addresses a key gap in the literature: it moves beyond average or individualised treatment-effect estimation to reveal subgroups of treatment response.
+Our empirical results on both synthetic and real-world datasets demonstrate that the proposed method successfully characterises latent subgroups with divergent treatment effects.
+While we acknowledge that our model relies on empirically unverifiable assumptions, as is often the case in causal inference, our focus on subgroup detection for hypothesis generation mitigates this concern.
+Indeed, such subgroup identification serves as a starting point for more targeted RCTs, where further validation can (i) confirm the heterogeneity in responses, (ii) explore alternative treatment strategies in less responsive subgroups, and (iii) shape evidence-based clinical guidelines.
+In doing so, this work provides a flexible and practical tool for designing future RCTs.
+By informing both the selection of patient populations most likely to benefit and avoiding treatments with limited or harmful effects, our framework offers a meaningful hypothesis-generating tool for clinicians, policymakers, and researchers striving to optimise patient care and resource allocation.
+Acknowledgements
+The authors acknowledge the partial support of the UKRI Medical Research Council (programme numbers MC_UU_00002/5 and MC_UU_00002/2 and theme number MC_UU_00040/02 – Precision Medicine).
+VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
+FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z).
+Appendix A Related work
 While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a, b; Jeanselme et al., 2022) to identify different phenotypes of patients, the ML literature on phenotyping treatment effects with time-to-event outcomes remains sparse.
-The current focus is on estimating population or individual treatment effects in the static (Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020) and survival settings (Curth et al., 2021).
-While these approaches help to understand the population or individualised treatment effects, they do not identify subgroups that may benefit or be harmed by treatment.
+The current focus is on estimating population or individual treatment effects in the static (Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020) and survival settings (Katzman et al., 2018; Curth et al., 2021).
+While these approaches estimate ATE, CATE, or even ITE, they do not directly identify subgroups that may benefit or be harmed by treatment.
 Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
@@
 (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits.
-Athey & Imbens (2015, 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation.
-Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation.
-Each obtained split in the decision tree delineates two subgroups of treatment effect (Lipkovich et al., 2011; Loh et al., 2015).
+Athey and Imbens (2015, 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation.
+Wager and Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation.
+Each obtained split in the decision tree delineates two subgroups of treatment effect Lipkovich et al.
+(2011); Loh et al.
+(2015).
 Alternatively, McFowland III et al.
-(2018) propose pattern detection and Wang & Rudin (2022), causal rule set learning to uncover these subgroups.
-However, all these approaches rely on a local optimisation criterion (Lipkovich & Dmitrienko, 2014) and greedy split exploration.
+(2018) propose pattern detection and Wang and Rudin (2022), causal rule set learning to uncover these subgroups.
+However, all these approaches rely on a local optimisation criterion (Lipkovich and Dmitrienko, 2014) and greedy split exploration.
 Recently, Nagpal et al.
 (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
+Our work proposes to address these limitations through a joint optimisation, while avoiding parametric treatment response and proportional hazard assumptions.
 Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with survival outcomes that our work explores.
@@
 Similarly, Hu et al.
-(2021) propose Bayesian additive models and Zhu & Gallego (2020) propose a step-wise approach with propensity weighting to study observational data.
-Closest to our work, (Jia et al., 2021; Nagpal et al., 2023) propose to uncover subgroups within RCTs with survival outcomes.
+(2021) propose Bayesian additive models and Zhu and Gallego (2020) propose a step-wise approach with propensity weighting to study observational data.
+Closest to our work, (An et al., 2023; Jia et al., 2021; Nagpal et al., 2023; Perrin et al., 2024) propose to uncover subgroups within RCTs with survival outcomes.
+Perrin et al.
+(2024) compares different recursive and tree-based appraoches to perform this task in RCTs.
+An et al.
+(2023) introduces an expectation-maximisation approach to jointly fit a logistic model for subgroup attribution and a Cox model for survival estimation.
 Jia et al.
@@
 An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect.
-As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
-3 Method
-3.1 Problem setup
-Our goal is to uncover a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 4.
-with distinct treatment responses, guided by the observed times of occurrence of an outcome of interest.
-To this end, our model assigns patients to latent subgroups based on their covariates at the treatment decision time.
-For each subgroup, we estimate the probability of not observing the event of interest given time under treatment and under control regimes.
-These two distributions are known as survival functions, and their difference corresponds to the estimated treatment effect for a given group.
-Formally, consider the random variables associated with observed covariates $X$, the assigned binary treatment $A$, and the observed event time $T^{\prime}$.
-Following the potential outcomes formulation, $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under consistency (Assumption 1) where $T^{\prime}_{1}$ is the potential event time under treatment and $T^{\prime}_{0}$, under the control regime.
-Assumption 1 (Consistency).
-A patient’s observed event time is the potential event time associated with the observed treatment.
-Formally, this means $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the control and treatment regimes respectively.
-Central to our problem is the latent, unobserved, subgroup membership $Z$.
-As formalised in Assumption 2, we assume that the variable $Z$ mitigates the dependence between the covariates $X$ and the observed event time $T^{\prime}$.
-Assumption 2 (Mixture mitigation).
-The event time $T^{\prime}$ is determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
-In this setting, our interest lies in estimating the Subgroup Average Treatment Effect (SATE) for group $k$:
-$\tau_{k}(t)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,Z=k)$
-$-\mathbb{P}(T^{\prime}\geq t\mid A=0,Z=k)$ (1)
-This focus differs from the literature interest in estimating Individualised Treatment Effects (ITE), defined as follows (see App. A.1 for derivation from the potential outcomes formalisation):
-$\tau(t,x)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)$
-$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$
-$=\sum_{k}\tau_{k}(t)\mathbb{P}(Z=k\mid X=x)$ (Under Asm. 2)
-Estimating the previous quantities requires the accurate modelling of the survival distributions under the two treatment regimes.
-The central challenge in this estimation is that the counterfactual survival outcome is unobserved: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
-The absence of treatment randomisation in observational studies, e.g.
-clinicians might recommend more aggressive treatment to patients in more severe conditions, results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021).
-This prohibits the estimation of the survival functions through the maximisation of the observed outcomes likelihood.
-Under Assumptions 3 and 4, remedies such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity between treatment regimes (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Hernán & Robins, 2006; Shalit et al., 2017) allows to estimate both factual and counterfactual likelihood from the observed data by penalising the difference between the treated and non-treated populations.
-In our work, we adopt an Inverse Propensity Weighting (IPW) of the factual likelihood to estimate the overall likelihood, consisting in weighting each observation by the inverse probability of receiving treatment.
-Assumption 3 (Ignorability).
-The potential event times are independent of the treatment given the observed covariates, i.e.
-$A\perp\!\!\!\perp(T^{\prime}_{0},T^{\prime}_{1})\mid X$.
-Equivalently, no unobserved confounders impact both treatment and event time.
-Assumption 4 (Overlap / Positivity).
-Each patient has a non-zero probability of receiving the treatment, i.e. $\mathbb{P}(A\mid X)\in(0,1)$ where $(0,1)$ is the open interval, resulting in a non-deterministic treatment assignment.
-To express the factual likelihood, one needs to tackle a final challenge in time-to-event data: censoring.
-Patients may not observe the outcome of interest during the study period or exit the study for reasons independent of the outcome of interest.
-Formally, instead of observing $T^{\prime}$, one observes an indicator identifying whether the event of interest was observed $D$, and the associated observed time of event $T$, with $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is the random variable of the (right)-censoring time.
-When a patient is censored, $T=C$ and $D=0$; otherwise, the event under treatment regime $A$ is observed and $T=T^{\prime}_{A}$ and $D=1$.
-As ignoring the censoring process biases the likelihood, we assume, as commonly done in survival analysis, that this process is uninformative, formalised as:
-Assumption 5 (Non-informative censoring).
-The censoring time $C$ is independent of the time of the event of interest $T^{\prime}$ given the covariates $X$. Formally, $T^{\prime}\perp\!\!\!\perp C\mid X$.
-Under the previous assumptions, we can maximise the overall likelihood by optimising for the following weighted factual log-likelihood $l_{F}$:
-$l_{F}=$ $\sum_{i,d_{i}=1}w_{i}\log\left(\frac{\partial S(t\mid x_{i},a_{i})}{\partial t}\bigg{|}_{t=t_{i}}\right)$
-$+\sum_{i,d_{i}=0}w_{i}\log S(t_{i}\mid x_{i},a_{i})$ (2)
-where $i$ is the patient index, $(x_{i},t_{i},a_{i},d_{i})$ are realisations of the associated variables $X,T,A$ and $D$, and $w_{i}$ is the inverse propensity weighting correction for patient $i$.
-Under Assumption 2, the survival corresponds to the weighted survival over the different subgroups:
-$S(t\mid x,a):=\mathbb{P}(T^{\prime}\geq t\mid x,a)$
-$=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid x,a)\mathbb{P}(T^{\prime}\geq t\mid Z=k,a)$ (3)
-In summary, Figure 2 describes all variables and dependences assumed in the studied problem with a directed acyclic graph.
-Given the observed $X$, $A$, $T$ and $D$, our aim is to estimate the latent structure $Z$ and the associated survival distributions under the two potential treatment regimes $T^{\prime}_{0}$ and $T^{\prime}_{1}$.
-Figure 2: Graphical representation between covariates ($X$), treatment ($A$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
-Figure 3: Causal Survival Clustering (CSC) architecture.
-Latent parameter $u_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the survival under both treatment regimes.
-$G$ assigns the probability of belonging to each subgroup given the patient’s covariate(s) $x$.
-To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood.
-In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
-3.2 Estimating the quantities of interest
-The previous section discussed the quantities one must estimate —here parametrised by neural networks [^2] When interpretability is a key concern, one can consider a linear model for both subgroup assignment and IPW components, as explored in Appendix B.4.— to uncover subgroups of treatment effects: the assignment function $\mathbb{P}(Z\mid X)$, the distributions characterising survival under the two treatment regimes $\mathbb{P}(T^{\prime}\geq t\mid A,Z=k)$, and the IPW weights $w$.
-Figure 3 illustrates the overall architecture and the neural networks used to estimate these quantities.
-Subgroup assignment.
-Similar to (Jeanselme et al., 2022), a multi-layer perceptron $G$ with a final Softmax layer assigns each patient characterised by covariates $x$ its probability of belonging to each subgroup, characterised through a $K$-dimensional vector of probabilities.
-$G(x):=[\mathbb{P}(Z=k\mid x)]_{k=1}^{K}$
-Survival distributions.
-Each subgroup $k$ is represented by a parameter vector $u_{k}\in\mathbbm{R}^{L}$ of dimension $L$, a latent parametrisation of the survival functions learnt through backpropagation.
-The vector $u_{k}$ is concatenated with $t$ and used as input to a neural network $M$ with monotonic outcomes [^3] To ensure this constraint, we apply a square function on all weights as proposed in (Jeanselme et al., 2022, 2023).
-and a final SoftPlus layer to ensure positivity.
-We train this neural network to model survival through the following transformation — ensuring that no probability is assigned to negative times, a limitation raised concerning previous monotonic neural networks (Shchur et al., 2020):
-$\mathbb{P}(T^{\prime}_{A=a}\geq t\mid Z=k):=\exp(-t\cdot M(u_{k},t)[a])$
-Inverse propensity weighting.
-Under treatment randomisation, as in RCTs, the previous components would accurately model the survival outcomes under both treatment regimes and identify subgroups of treatment effects by maximising the factual likelihood.
-As previously discussed, to account for the treatment non-randomisation in observational studies, we weight the factual likelihood using the propensity of receiving treatment estimated through a multi-layer perceptron $W$ with a final sigmoid transformation as:
-$W(x):=\mathbb{P}(A=1\mid x)$
-From this estimated probability, we further truncate the propensity score (Austin & Stuart, 2015) to avoid extreme values for the estimated weights, which would result in unstable treatment effect estimates.
-In this context, the weights $w_{i}$ are defined as:
-$\forall i,w_{i}^{\prime}:=$ $a_{i}W(x_{i})+(1-a_{i})(1-W(x_{i}))$
-$w_{i}:=$ $\max(0.05,\min(w_{i}^{\prime},0.95))$ (4)
-3.3 Training procedure
-First, the model $W$ is trained to predict the binary treatment assignment probability by minimising the cross-entropy of receiving treatment.
-Then, training all other components relies on maximising the weighted log-likelihood introduced in Equation (2).
-The use of monotonic neural networks results in the efficient and exact computation of the log-likelihood as a forward path provides the estimated survival and automatic differentiation of the monotonic neural networks’ outcomes readily provides the derivative necessary for computing the log-likelihood (Jeanselme et al., 2022, 2023; Rindt et al., 2022).
-4 Experimental analysis
-As counterfactuals are unknown in observational data, we adopt—as is common practice in this research field—a synthetic dataset in which underlying survival distributions and group structure are known.
-As a real-world case study, Appendix C accompanies these synthetic results with the analysis of heterogeneity in adjuvant radiotherapy responses for patients diagnosed with breast cancer.
-Our code to produce the synthetic dataset, the model, and all experimental results is available on Github [^4] https://github.com/Jeanselme/CausalNeuralSurvivalClustering.
-4.1 Data generation
-We generate a population of $3,000$ equally divided into $K=3$ subgroups.
-We draw 10 covariates from normal distributions with different centres and survival distributions for each treatment regime following Gompertz distributions (Pollard & Valkovics, 1992) parametrised by group membership and individual covariates.
-Note that this setting relaxes the independence of outcomes and covariates assumed by the proposed model (Assumption 2).
-This choice reflects a setting in which survival distributions under treatment or control regimes are different functions of the covariates.
-This simulation is more likely to capture the complexity of real-world responses, in contrast to traditional evaluations of subgroup analysis, which often assume a linear treatment response.
-Further, we implement two treatment assignment scenarios: a randomised assignment in which treatment is independent of patient covariates, similar to RCTs (Randomised), and one in which treatment is a function of the patient covariates as in observational studies (Observational).
-Finally, non-informative censoring times are drawn following a Gompertz distribution.
-Details of the generative process of our synthetic dataset are deferred to Appendix B.1.
-Further, Appendix B.5 explores alternative datasets to demonstrate the flexibility of the proposed methodology.
-4.2 Empirical settings
-Benchmark methods.
-We compare the proposed approach (CSC) against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE (Nagpal et al., 2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups.
-This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted.
-The central differences with our proposed approach are that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response.
-This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival.
-Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses.
-By contrast, our approach identifies treatment subgroups while considering survival without constraining the treatment response.
-We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, compared to a group with the same treatment responses but with better survival odds.
-Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
-For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups.
-Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
-Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
-Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately.
-First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in (Nagpal et al., 2022a), referred to as Kmeans + TE in the following.
-Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data, and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters.
-Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them.
-As proposed in the literature, we use a survival tree to estimate response under each treatment regime and then use a KMeans (MacQueen et al., 1967) on the estimated treatment effects, computed as the difference between survival estimates.
-Note that these methodologies assume treatment randomisation and the clustering is uninformed by the covariates.
-For details on training and hyperparameter optimisation, we refer to App.
-B.2.
-Evaluation.
-In the synthetic experiments, the subgroup structure is known.
-We measure the adjusted [^6] Random patient assignment results in an adjusted Rand-Index of 0.
-Rand-Index (Rand, 1971), which quantifies how the estimated assignment aligns with the known underlying group structure.
-Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
-[…]
-where $\hat{\tau}_{\hat{k}}$ is the estimated treatment effect for subgroup $\hat{k}=\operatorname*{arg\,max}_{l}\mathbb{E}_{x\in k}(Z=l\mid x)$, i.e.
-the most likely assigned cluster for patients in the underlying $k$ cluster, and $\tau_{k}$ is the ground truth.
-4.3 Treatment effect recovery
-Figure 4: Averaged negative log-likelihood across 5-fold cross-validation given the number of subgroups $K$ under the ”Observational” treatment assignment with the shaded area representing 95% CI.
-The log-likelihood presents an elbow around the underlying number of subgroups.
-Recovering the underlying number of subgroups.
-For all methodologies, we must choose the number of subgroups a priori.
-An important question is, therefore, whether we can identify the true number of subgroups with our model.
-Figure 4 presents the average negative log-likelihood obtained by cross-validation for models with different numbers of subgroups $K$.
-The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood.
-Using this heuristic, the optimal choice for $K$ is $3$, which aligns with the underlying generative process.
-App.
-B.3 shows that when the number of clusters is misspecified, the estimated treatment effects are unstable, the variance associated with estimated groups is hence an additional tool to select the number of clusters.
-This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which separate clustering from treatment effect estimates.
-In these methods, survival outcomes cannot directly guide the choice of $K$, whereas it is directly reflected in the proposed methodology’s performance.
-Discovering subgroups.
-Next, we present the quantitative results of our method against the benchmark methods.
-Table 1 summarises the performance of the different methodologies under the two studied scenarios.
-Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival clusters.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.807 (0.055) 0.037 (0.008) 0.023 (0.008) 0.015 (0.003) CSC Unadjusted 0.802 (0.050) 0.039 (0.011) 0.021 (0.009) 0.014 (0.003) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.255 (0.159) - 0.084 (0.005) - CMHE ($K=L$) 0.345 (0.104) 0.243 (0.042) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
-Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation.
-Best performance per column and simulation scenario are marked in bold, second best in italic.
-’-’ describes when the methodology diverges.
-Our proposed CSC method and its unadjusted variant best recover the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
-CSC outperforms all CMHE alternatives, the current state-of-the-art method.
-CMHE’s parametrisation and assumptions explain this difference.
-Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log hazards.
-Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data.
-In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes.
-Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods.
-These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
-CSC presents the best performance in identifying the underlying subgroup.
-Our proposed CSC method outperforms all approaches on the Rand-Index.
-In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are unaligned.
-KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index.
-Consequently, the identified clusters differ in their treatment responses, as shown by the large IAE.
-The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE.
-However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates.
-This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
-Treatment assignment correction improves subgroup identification in observational settings.
-The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation.
-The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context.
-However, CSC better recovers the different groups as shown by the Rand-Index.
-Finally, Appendices B.5.1 to B.5.4 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
-Real-world heterogeneity analysis in adjuvant radiotherapy responses (App. C).
-Using the Surveillance, Epidemiology, and End Results program(Seer) [^7] Available at https://seer.cancer.gov/, App.
-C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research.
-Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment.
-Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis.
-Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
-5 Conclusion
-Understanding heterogeneity in treatment effect is essential for guiding clinical practice and informing the development of new treatments.
-Motivated principally by how real-world clinical guidelines depend on patient subgroups, this work introduces a neural network-based framework for identifying treatment-response subgroups in observational time-to-event data.
-By leveraging large-scale observational datasets and survival analysis with causal inference methods, our approach addresses a key gap in the literature: it moves beyond average or individualised treatment-effect estimation to reveal subgroups of treatment response.
-Our empirical results on both synthetic and real-world datasets demonstrate that the proposed method successfully characterises latent subgroups with divergent treatment effects.
-While we acknowledge that our model depends on empirically unverifiable assumptions as often in causal inference, our focus on subgroup detection for hypothesis generation lessens this concern.
-Indeed, such subgroup identification serves as a starting point for more targeted RCTs, where further validation can (i) confirm the heterogeneity in responses, (ii) explore alternative treatment strategies in less responsive subgroups, and (iii) shape evidence-based clinical guidelines.
-In doing so, this work contributes a flexible and practical tool for the design of future RCTs.
-By informing both the selection of patient populations most likely to benefit and avoiding treatments with limited or harmful effects, our framework offers a meaningful hypothesis-generating tool for clinicians, policymakers, and researchers striving to optimise patient care and resource allocation.
-Acknowledgements
-The authors acknowledge the partial support of the UKRI Medical Research Council (programme numbers MC_UU_00002/5 and MC_UU_00002/2 and theme number MC_UU_00040/02 – Precision Medicine).
-VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.
-FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z).
-Appendix A Proof
-A.1 Individualised Treatment Effect
-This section derives the individualised treatment effect expression introduced in Section 3.
+As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational settings.
+Appendix B Proof
+This section derives the conditional average treatment effect expression introduced in Sec. 2.
 $\tau(t,x):=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)-\mathbbm{1}(T_{0}\geq t)\mid X=x)$
-$=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)\mid X=x)$
-$-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
-$=\mathbb{P}($ $T^{\prime}\geq t\mid A=1,X=x)$
-$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under assumption 1 and 3)
-Appendix B Synthetic analysis
-B.1 Data generation
+$=\mathbb{E}($ $\mathbbm{1}(T_{1}\geq t)\mid X=x)-\mathbb{E}(\mathbbm{1}(T_{0}\geq t)\mid X=x)$
+$=\mathbb{P}($ $T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (Under Asmp. 1 and 3)
+Appendix C Synthetic analysis
+C.1 Data generation
 We consider a synthetic population of $N=3,000$ patients with 10 associated covariates $X\in\mathbb{R}^{10}$ divided into $K=3$ subgroups.
-The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to (Nagpal et al., 2022a).
+The following data generation does not aim to mimic a particular real-world setting but follows a similar approach to Nagpal et al.
+(2022a).
 The following describes our generation process:
@@
 Each patient’s membership $Z$ is drawn from a multinomial with equal probability.
-Group membership informs the two first covariates through the parametrisation of the bivariate normal distribution with centres $c_{k}$ equal to $(0,2.25)$, $(-2.25,-1)$, and $(2.25,-1)$.
+Group membership informs the first two covariates through the parametrisation of the bivariate normal distribution with centres $c_{k}$ equal to $(0,2.25)$, (2.25,1)(-2.25,-1), and (2.25,1)(2.25,-1).
 All other covariates are drawn from standard normal distributions.
@@
 with MVN denoting a multivariate normal distribution, $c^{\prime}_{k}$ random cluster centres, and $I^{n}$, the identity covariance matrix of dimension $n$.
-Note that we introduce $Y$ to present a covariates structure that is independent of the treatment responses.
-(a) $K=2$
-(b) $K=3$
-(c) $K=4$
-Figure 5: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross validated average estimated treatment effect. Lines in grey corresponds to the ground truth.
+Note that we introduce $Y$ to present a covariate structure that is independent of the treatment responses.
+[…]
+[…]
+Figure 6: Sensitivity to the number of clusters used to train CSC. Lines in blue represent the cross-validated average estimated treatment effect. Lines in grey correspond to the ground truth.
 Treatment response.
@@
 Assuming a treatment assignment probability of 50%, we assign each patient to a given treatment.
-We propose two treatment assignment strategies reflecting a RCT and an observational setting, denoted as Randomised and Observational.
+We propose two treatment assignment strategies reflecting an RCT and an observational setting, denoted as Randomised and Observational.
 Randomised consists of a Bernoulli draw using the realisation of $P$.
@@
 Our goal is to model the treatment effect and identify the underlying subgroup structure $Z$ from the observed $X,T,D,A$ with $T\mid A=a,T_{0},T_{1},C=t_{0},t_{1},c:=\min(c,(1-a)\times t_{0}+a\times t_{1})$ and $D=\mathbbm{1}_{C>T}$.
-B.2 Training and hyperparameter optimisation
+C.2 Training and hyperparameter optimisation
 Training.
@@
 For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search.
-All models were optimised for 1000 epochs using an Adam optimiser (Kingma & Ba, 2015) with early stopping.
+All models were optimised for 1000 epochs using an Adam optimiser (Kingma and Ba, 2015) with early stopping.
 Hyperparameter optimisation.
@@
 Tree-based approach to estimate treatment effect had a depth limited to: 3, 5 or unconstrained, and the minimum sample split was chosen between 6, 12 or 24.
-B.3 Sensitivity to choice of K
-In Section 4, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$.
-While this outcome-guided selection of the number of clusters is a core strength of the proposed methodology in comparison to existing strategies, we explore how the identified treatment effects change under a misspecified model which estimates 2 and 4 clusters.
+C.3 Sensitivity to choice of K
+In Sec.
+3, we describe how the proposed methodology presents an elbow in the likelihood as a function of the number of clusters around the underlying number of clusters $K=3$.
+While this outcome-guided selection of the number of clusters is a core strength of the proposed methodology in comparison to existing strategies, we explore how the identified treatment effects change under a misspecified model that estimates 2 and 4 clusters.
 This analysis examines the risk associated with misspecifying the number of clusters.
@@
 CSC 0.797 (0.043) 0.042 (0.011) 0.051 (0.026) 0.022 (0.004) CSC Unadjusted 0.742 (0.073) 0.047 (0.019) 0.030 (0.029) 0.020 (0.003) CSC Linear 0.819 (0.041) 0.047 (0.015) 0.031 (0.022) 0.022 (0.003)
-Table 2: Cross-validated performance (with standard deviation in parenthesis).
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.643 (0.067) 0.060 (0.016) 0.021 (0.007) 0.070 (0.013) CSC Unadjusted 0.643 (0.071) 0.064 (0.013) 0.018 (0.008) 0.060 (0.010) CMHE ($K=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($L=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($K=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.788 (0.113) 0.040 (0.017) 0.041 (0.017) 0.086 (0.043) CSC Unadjusted 0.626 (0.170) 0.080 (0.015) 0.031 (0.009) 0.095 (0.049) CMHE ($K=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($L=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($K=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
-Table 3: Cross-validated performance (with standard deviation in parenthesis) when clusters are of different sizes.
-Figure 5 illustrates the cross-validated subgroup treatment effects when the model is trained with $K=2,3$ and $4$ clusters, despite the existence of 3 clusters.
-When using a number of clusters different than indicated by the elbow heuristic, the identified clusters are less stable as not capturing the true underlying treatment responses, as shown by the increased standard deviation.
-Uncertainty increases with misspecified models, providing an additional tool to select the appropriate number of cluster: the stability of the estimated clusters over cross-validation indicates a better alignment with underlying distribution.
-B.4 Linear Modelling
-While the flexibility of neural network allows flexibility of the survival distribution, one could consider more interpretable assignment $G$ and treatment $W$ components than the neural network approach proposed in the main text.
+Table 3: Cross-validated performance (with standard deviation in parentheses).
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.643 (0.067) 0.060 (0.016) 0.021 (0.007) 0.070 (0.013) CSC Unadjusted 0.643 (0.071) 0.064 (0.013) 0.018 (0.008) 0.060 (0.010) CMHE ($L=3$) 0.589 (0.019) 0.179 (0.011) 0.096 (0.008) 0.086 (0.004) CMHE ($M=3$) 0.072 (0.051) 0.208 (0.004) 0.211 (0.027) 0.122 (0.011) CMHE ($M=L$) 0.471 (0.022) 0.241 (0.011) 0.264 (0.016) 0.173 (0.010) KMeans + TE 0.336 (0.010) 0.078 (0.008) 0.029 (0.007) 0.237 (0.017) Virtual Twins 0.541 (0.163) 0.033 (0.025) 0.052 (0.018) 0.087 (0.051) Observational CSC 0.788 (0.113) 0.040 (0.017) 0.041 (0.017) 0.086 (0.043) CSC Unadjusted 0.626 (0.170) 0.080 (0.015) 0.031 (0.009) 0.095 (0.049) CMHE ($L=3$) 0.611 (0.022) 0.179 (0.014) 0.103 (0.005) 0.089 (0.005) CMHE ($M=3$) 0.084 (0.034) 0.211 (0.007) 0.209 (0.011) 0.127 (0.009) CMHE ($M=L$) 0.462 (0.062) 0.252 (0.012) 0.286 (0.008) 0.096 (0.010) KMeans + TE 0.341 (0.020) 0.045 (0.009) 0.043 (0.009) 0.275 (0.017) Virtual Twins 0.534 (0.166) 0.040 (0.015) 0.051 (0.024) 0.084 (0.035)
+Table 4: Cross-validated performance (with standard deviation in parentheses) when clusters are of different sizes.
+Figure 6 illustrates the cross-validated subgroup treatment effects when the model is trained with $K=2,3$, and $4$ clusters, despite the existence of 3 clusters.
+When using a number of clusters different from that indicated by the elbow heuristic, the identified clusters are less stable, as they do not capture the true underlying treatment responses, as shown by the increased standard deviation.
+Uncertainty increases with misspecified models, providing an additional tool to select the appropriate number of clusters: the stability of the estimated clusters over cross-validation indicates a better alignment with the underlying distribution.
+C.4 Linear modelling
+While the flexibility of neural networks allows flexibility of the survival distribution, one could consider more interpretable assignment $G$ and treatment $W$ components than the neural network approach proposed in the main text.
 This section explores the recovery of treatment effects and subgroups when using linear regressions, i.e.
@@
 Table 3 summarises the results presented in the main paper with the additional linear alternative in which both the treatment and assignment networks are replaced by a linear model, referred as CSC Linear.
-These results demonstrate that enforcing linear relation between covariates and the different quantities of interest can reduce the risk of overfitting in small datasets as shown by the improved Rand-Index.
-Note that under larger dataset or more complex relation between covariates, group membership and treatment, one should consider more flexible modelling, such as the proposed neural approach.
-B.5 Alternative data generations
+These results demonstrate that enforcing a linear relation between covariates and the different quantities of interest can reduce the risk of overfitting in small datasets, as shown by the improved Rand-Index.
+Note that under a larger dataset or a more complex relation between covariates, group membership, and treatment, one should consider more flexible modelling, such as the proposed neural approach.
+C.5 Alternative data generations
 In this section, we explore alternative data-generative processes to demonstrate the robustness of the proposed strategy under different settings.
-B.5.1 Unequal cluster size
-Section B.1 presents an equally-distributed population over the different clusters.
+C.5.1 Unequal cluster size
+Sec.
+C.1 presents an equally distributed population over the different clusters.
 In medical settings, the underlying subgroups may differ in size.
-This section explores an alternative scenario in which the population is distributed over the three clusters as follows: 62.5%, 25% and 12.5%.
+This section explores an alternative scenario in which the population is distributed over the three clusters as follows: 62.5%, 25%, and 12.5%.
 Table 4 summarises the performance in this simulation, echoing the main text’s conclusions.
 The proposed method best recovers the different subgroups and presents one of the best estimates of subgroups’ treatment effects.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.310 (0.187) 0.236 (0.103) 0.110 (0.033) 0.149 (0.106) CSC Unadjusted 0.443 (0.285) 0.176 (0.098) 0.087 (0.046) 0.132 (0.112) CMHE ($K=3$) 0.113 (0.100) 0.296 (0.040) 0.137 (0.042) 0.188 (0.080) CMHE ($L=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($K=L$) 0.079 (0.119) 0.359 (0.068) 0.182 (0.089) 0.240 (0.058) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.764 (0.015) 0.027 (0.008) 0.011 (0.001) 0.047 (0.008) CSC Unadjusted 0.729 (0.022) 0.033 (0.008) 0.010 (0.003) 0.039 (0.009) CMHE ($K=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($L=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($K=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
-Table 4: Cross-validated performance (with standard deviation in parenthesis) with varying $N$ under observational treatment setting.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.830 (0.047) 0.051 (0.008) 0.023 (0.013) 0.020 (0.005) CSC Unadjusted 0.857 (0.033) 0.045 (0.006) 0.030 (0.015) 0.015 (0.003) CMHE ($K=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.134 (0.090) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.718 (0.198) 0.061 (0.050) 0.057 (0.048) 0.023 (0.005) CSC Unadjusted 0.804 (0.045) 0.037 (0.006) 0.023 (0.008) 0.025 (0.007) CMHE ($K=3$) 0.385 (0.039) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($L=3$) 0.318 (0.140) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($K=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012)
-Table 5: Cross-validated performance (with standard deviation in parenthesis) with varying treatment rates under observational settings.
-B.5.2 Number of points
-Section B.1 describes a data generation with $3,000$ patients.
-This section presents results when there are $300$ and $30,000$ patients.
+C.5.2 Number of points
+Sec.
+C.1 describes a data generation with $3,000$ patients.
+This section presents results for $300$ and $30,000$ patients.
 A smaller number of patients may impact the neural network capacity to identify underlying subgroups of treatment effect.
 A larger number makes non-randomisation less of a concern.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 300 CSC 0.310 (0.187) 0.236 (0.103) 0.110 (0.033) 0.149 (0.106) CSC Unadjusted 0.443 (0.285) 0.176 (0.098) 0.087 (0.046) 0.132 (0.112) CMHE ($L=3$) 0.113 (0.100) 0.296 (0.040) 0.137 (0.042) 0.188 (0.080) CMHE ($M=3$) 0.131 (0.067) - - 0.254 (0.022) CMHE ($M=L$) 0.079 (0.119) 0.359 (0.068) 0.182 (0.089) 0.240 (0.058) KMeans + TE 0.050 (0.110) 0.251 (0.072) 0.154 (0.065) 0.267 (0.049) Virtual Twins 0.296 (0.115) 0.159 (0.072) 0.172 (0.067) 0.101 (0.043) 30,000 CSC 0.764 (0.015) 0.027 (0.008) 0.011 (0.001) 0.047 (0.008) CSC Unadjusted 0.729 (0.022) 0.033 (0.008) 0.010 (0.003) 0.039 (0.009) CMHE ($L=3$) 0.567 (0.159) 0.094 (0.011) 0.059 (0.018) 0.142 (0.034) CMHE ($M=3$) 0.622 (0.040) 0.078 (0.001) 0.125 (0.003) 0.226 (0.003) CMHE ($M=L$) 0.638 (0.008) 0.059 (0.001) 0.220 (0.007) 0.140 (0.007) KMeans + TE 0.000 (0.000) 0.087 (0.004) 0.147 (0.008) 0.208 (0.004) Virtual Twins 0.650 (0.030) 0.011 (0.003) 0.009 (0.000) 0.010 (0.004)
+Table 5: Cross-validated performance (with standard deviation in parentheses) with varying NN under an observational treatment setting.
 Table 5 presents the performance with these different population sizes under an observational treatment assignment.
-A first observation is that all methodologies present better performance with larger number of points.
-Further, baselines that do not account for the assignment mechanisms present lower performance with $N=300$ as non-randomisation has an increased impact on treatment effect estimate with a smaller population.
+A first observation is that all methodologies present better performance with a larger number of points.
+Further, baselines that do not account for the assignment mechanisms present lower performance with $N=300$ as non-randomisation has an increased impact on treatment effect estimates with a smaller population.
 This difference decreases when $N=30,000$ with the virtual twins approach presenting the best recovery of the treatment effects.
-However, throughout the different settings, the proposed CSC presents the best Rand-Index indicating a good recovery of the underlying structure.
-B.5.3 Impact of treatment rate
-Section B.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.
-Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.559 (0.114) 0.024 (0.009) 0.030 (0.017) 0.030 (0.006) CSC Unadjusted 0.481 (0.069) 0.026 (0.005) 0.041 (0.026) 0.032 (0.009) CMHE ($K=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($L=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($K=L$) 0.254 (0.138) 0.074 (0.027) 0.052 (0.021) 0.149 (0.018) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.584 (0.151) 0.015 (0.007) 0.028 (0.024) 0.035 (0.003) CSC Unadjusted 0.556 (0.249) 0.023 (0.008) 0.051 (0.025) 0.040 (0.006) CMHE ($K=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($L=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($K=L$) 0.293 (0.125) 0.063 (0.012) 0.051 (0.017) 0.140 (0.011) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
-Table 6: Cross-validated performance (with standard deviation in parenthesis) when all covariates clusters presents different treatment responses.
-This setting aligns with the assumptions made by the step-wise approaches.
+However, throughout the different settings, the proposed CSC presents the best Rand-Index, indicating a good recovery of the underlying structure.
+C.5.3 Impact of treatment rate
+Sec. C.1 assumes 50% of the population receives treatment. This section explores when 25% and 75% of the population receive treatment under the non-randomised treatment setting.
 Table 6 presents the performance under these different treatment rates in an observational setting.
 These results highlight CSC’s capacity to identify the underlying subgroups with the highest Rand-Index and one of the best cluster treatment effect recovery in these settings.
-B.5.4 Aligned covariates and treatment effect structure
-Section B.1 presents a data generation with the last 8 covariates having a clustered structures independent of the treatment response.
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.830 (0.047) 0.051 (0.008) 0.023 (0.013) 0.020 (0.005) CSC Unadjusted 0.857 (0.033) 0.045 (0.006) 0.030 (0.015) 0.015 (0.003) CMHE ($L=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($M=3$) 0.134 (0.090) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($M=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.718 (0.198) 0.061 (0.050) 0.057 (0.048) 0.023 (0.005) CSC Unadjusted 0.804 (0.045) 0.037 (0.006) 0.023 (0.008) 0.025 (0.007) CMHE ($L=3$) 0.385 (0.039) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($M=3$) 0.318 (0.140) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($M=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012)
+Table 6: Cross-validated performance (with standard deviation in parentheses) with varying treatment rates under observational settings.
+C.5.4 Aligned covariates and treatment effect structure
+Sec.
+C.1 presents a data generation with the last 8 covariates having a clustered structure independent of the treatment response.
 This section explores a setting where all clusters in the covariates are associated with different treatment responses of interest.
@@
 $X_{[3-10]}\sim\text{MVN}(0,I^{8})$
+Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.559 (0.114) 0.024 (0.009) 0.030 (0.017) 0.030 (0.006) CSC Unadjusted 0.481 (0.069) 0.026 (0.005) 0.041 (0.026) 0.032 (0.009) CMHE ($L=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($M=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($M=L$) 0.254 (0.138) 0.074 (0.027) 0.052 (0.021) 0.149 (0.018) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.584 (0.151) 0.015 (0.007) 0.028 (0.024) 0.035 (0.003) CSC Unadjusted 0.556 (0.249) 0.023 (0.008) 0.051 (0.025) 0.040 (0.006) CMHE ($L=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($M=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($M=L$) 0.293 (0.125) 0.063 (0.012) 0.051 (0.017) 0.140 (0.011) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
+Table 7: Cross-validated performance (with standard deviation in parentheses) when all covariate clusters present different treatment responses.
+This setting aligns with the assumptions made by the step-wise approaches.
 Table 7 summarises the performance in this setting, evidencing an improvement of the KMeans+TE approach as this model makes this assumption.
@@
 Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses.
-This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across different and even adverserial settings.
-Appendix C Case study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
-To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017.
-Following (Lee et al., 2018; Danks & Yau, 2022; Jeanselme et al., 2023), we select women who died from the condition or from cardiovascular diseases.
-From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes.
-To this end, we subselect patients with recorded treatment and who received chemotherapy.
-These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  (Danks & Yau, 2022) for further description).
-Selection of $K$.
-From the selected population, our aim is to identify heterogeneous responses to adjuvant treatment.
-The first challenge is the selection of the number of groups to use ($K$).
-We advise to follow medical actionability and consider the change in treatment effects and size of the subgroups when increasing this parameter.
-In the absence of experts’ intuition, one may rely on an elbow rule heuristic over $l_{F}$.
-Using Figure 6, the negative log-likelihood presents an elbow for $K=2$.
+This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across different and even adversarial settings.
+Appendix D Additional results: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
+This section presents additional results on the heterogeneity of adjuvant radiotherapy responses in the Seer dataset.
+D.1 Selection of $K$
+Using Figure 7, one can use the elbow rule on the negative log-likelihood. Following this heuristic, our analysis uses $K=2$ subgroups.
 Figure 7: Cross-validated negative log-likelihood as a function of the number of groups ($K$).
-Treatment effect subgroups.
-In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy.
-This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist (Lazzari et al., 2023), making this setting more likely to meet the positivity assumption (Assumption 4), necessary to study causality in observational data.
-(a) Causal Survival Clustering
-(b) Cox Mixture of Heterogeneous Effects
+D.2 Factors impacting subgroup membership
+Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariate.
+Using a permutation test, we identify the covariates that most impact the likelihood associated with the proposed CSC model.
+Figure 8 displays the 10 covariates which most impact the likelihood, and therefore are associated with the different treatment response subgroups.
+This proposed analysis provides a tool to validate the model’s medical relevance by studying which covariates are associated with outcomes.
+D.3 Baselines’ identified subgroups
+[…]
+[…]
 Figure 9: Averaged treatment effect subgroups across 5-fold cross-validation observed in the Seer dataset with the shaded areas representing 95% CI.
-Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
-RMST at 5 years Population % Treated % Distant Lymph Nodes HER2 Positive ER Positive Subgroup 0 0.01 (0.00) 94.6% 55.6% 1.2 (5.90) 17.5% 46.5% Subgroup 1 0.84 (0.11) 5.4% 45.1% 20.5 (15.56) 23.4% 50.4%
-Table 7: Causal Survival Clustering subgroups’ characteristics in the Seer cohort described through percentage / mean (std).
-Uncovering treatment response.
-Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE.
-For this latter, the number of subgroups $K$ is selected through hyperparameter tuning, leading to $K=1$ in every fold.
-As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups.
-As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects.
+Our main analysis describes how to use the proposed methodology to study a medical application.
+In this section, we analyse the groups one would identify with the previously described baselines.
+Figure 9 displays the identified treatment effect subgroups using all considered baselines.
+For CMHE, the number of subgroups $K$ is selected through hyperparameter tuning, leading to $K=1$ in every fold.
+As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects compared to CMHE.
 First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions.
@@
 Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
-Using a permutation test, we identify as the covariates that most impact the likelihood associated with the model.
-Figure 8 displays the 10 covariates most indicative of the different treatment response subgroups.
-Table 7 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) (Royston & Parmar, 2013).
-Both methodologies identify a population with limited treatment response.
-However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
-The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy.
-However, our methodologies remain hypothesis-generating tools, requiring further experimental validation, particularly due to potential confounding through hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options.
-The quality of available covariates limits this analysis and only serves as an example to medical practitioners to identify subgroups of treatment responses from observational data.
+While both methodologies identify a group with limited response, our proposed methodology distinguishes a second group with improved treatment response.
+Additionally, the KMeans approach identifies clusters that resemble those found with CSC, though the detected positive response is comparatively smaller.
+In contrast, the virtual twins method identifies a subgroup that exhibits a larger treatment effect, while the majority shows a negative response to treatment.
+Importantly, both methods broadly corroborate our findings.
+However, CSC presents a key strength by jointly learning group membership and treatment effects, ensuring that the covariates defining each subgroup are directly associated with the outcome – a property that two-step approaches cannot guarantee.